### v1

# Double Fairness Policy Learning: Integrating Action Fairness and Outcome Fairness in Decision-making

Abstract
Fairness is a central pillar of trustworthy machine learning, especially in domains where accuracy- or profit-driven optimization is insufficient. While most fairness research focuses on supervised learning, fairness in policy learning remains less explored. Because policy learning is interventional, it induces two distinct fairness targets: action fairness (equitable action assignments) and outcome fairness (equitable downstream consequences). Crucially, equalizing actions does not generally equalize outcomes when groups face different constraints or respond differently to the same action. We propose a novel double fairness learning (DFL) framework that explicitly manages the trade-off among three objectives: action fairness, outcome fairness, and value maximization. We integrate fairness directly into a multi-objective optimization problem for policy learning and employ a lexicographic weighted Tchebyshev method that recovers Pareto solutions beyond convex settings, with theoretical guarantees on the regret bounds. Our framework is flexible and accommodates various commonly used fairness notions. Extensive simulations demonstrate improved performance relative to competing methods. In applications to a motor third-party liability insurance dataset and an entrepreneurship training dataset, DFL substantially improves both action and outcome fairness while incurring only a modest reduction in overall value.
Keywords: Fairness; Trustworthiness; Policy Learning; Multi-objective Optimization.
1 Introduction
As a core aspect of trustworthiness, fairness has become central to machine learning (Dwork et al., 2012; Zemel et al., 2013; Grgic-Hlaca et al., 2016; Hardt et al., 2016; Kusner et al., 2017; Wu et al., 2019; Tan et al., 2022; Chen et al., 2023), motivated by ethical concerns about the societal impacts of algorithmic decision-making. In many domains such as healthcare, insurance, lending, and public programs, algorithmic systems can affect large populations at scale. Purely accuracy- or profit-driven optimization can amplify disparities, imposing disproportionate burdens on vulnerable groups, and exposing organizations to legal and reputational risk. Fairness is therefore not merely a normative ideal. It is an operational requirement for responsible deployment and a critical pillar of ethical AI.
Most existing work on algorithmic fairness has focused on supervised learning. In contrast, fairness in policy learning, where the goal is to maximize expected reward or social welfare, has received comparatively limited attention (Nabi et al., 2019; Tan et al., 2022; Viviano and Bradic, 2023; Fang et al., 2023; Wang et al., 2025). A key distinction is that policy learning is inherently interventional: a policy selects an action $A$, and that action alters subsequent outcomes. This leads to two conceptually distinct fairness targets. Action fairness concerns whether the policy treats individuals similarly in its action choices across sensitive groups (e.g., equalized treatment assignments). Outcome fairness instead asks whether the consequences (e.g., the benefits, welfare, or other fairness-relevant outcomes) induced by the policy are equitably distributed across groups. Importantly, equalizing actions does not, by itself, equalize outcomes when groups face different constraints, environments, or downstream responses to the same action. Consistent with this distinction, our analysis of a Belgian motor liability insurance dataset (Dutang and Charpentier, 2019) show that standard policy optimization methods can induce substantial disparities in both action and outcome fairness (Table 1 in Section 6).
This “action–outcome” gap arises naturally in practice. For instance, a government job training program might admit participants equally (action fair), unequal downstream constraints such as transportation access, caregiving burdens, or financial instability can still prevent some participants from completing the program or benefiting from it to the same extent. This could result in an inequitable final outcome (e.g., different job placement rates) even though the admission decision appears fair. In our analysis of such a training program in Section 6.2, we found a stark example of this disparity: one demographic group had a nearly 20% lower probability of securing an internship than another. This underscores the importance of outcome-level fairness in decision-making.
This work is also motivated by growing concerns about fairness in the insurance industry. Machine learning algorithms are increasingly used to automate risk assessment, pricing, and claims-related decisions, making trustworthy decision rules especially critical. For example, UnitedHealth has used the nH Predict platform (Talia, 2024) to assess patient risk and forecast outcomes in ways that inform coverage and pricing. While such tools can improve efficiency, they also raise concerns about algorithmic bias and discrimination against vulnerable populations. A recent class-action lawsuit alleges that the nH Predict algorithm was used to unlawfully deny or curtail rehabilitation care for seriously ill and elderly. [^1] See the news report at Healthcare Finance News.
The insurance context further illustrates why action fairness alone is it insufficient. In automobile insurance, a typical “action” is to classify a customer as low-risk vs. high-risk, which directly affects premiums. Action fairness focuses on whether similarly situated individuals across groups receive comparable risk labels and charged comparable premiums. However, even under perfect action fairness, two individuals with similar profiles receiving the same label and premium, their realized protection can still differ across groups. For example, claims may be approved at different rates, reimbursed differently, or processed with different delays across groups due to differences in documentation burdens, access to provider networks, or other downstream mechanisms. In such cases, the action is equal, but the policy-induced outcome, the expected net benefit or welfare a customer receives, remains unequal. Outcome fairness is designed to detect and mitigate precisely these disparities.
These findings highlight the need to balance the decision maker’s value objectives with fair treatment across different groups. Motivated by these concerns, we propose a novel double fairness policy learning framework that jointly targets (a) action fairness; (b) outcome fairness; and (c) maximization of the overall value function. We refer to this integrated goal as double fairness learning (DFL).
The DFL framework raises three central challenges. First, outcome fairness can only be controlled indirectly through the policy’s actions and the environment’s response: unlike action fairness, one cannot directly “intervene” on outcomes, so it is unclear when an outcome-fair (or doubly fair) policy exists at all. Second, the three objectives, action fairness, outcome fairness and value maximization, typically conflict, so it is rare for a single policy to optimize all three simultaneously. Hence, a principled framework is needed to characterize and manage these trade-offs, for which we adopt the concept of Pareto optimality. Third, computing the relevant Pareto fairness set is nontrivial in practice: standard linear scalarization approaches can miss Pareto-optimal solutions when the policy class is non-convex, as is typical for many policy parameterizations (Schulman et al., 2015).
We address these challenges with a unified methodology and theory. We first provide a systematic analysis of when outcome-fair and doubly fair policies can be achieved, yielding interpretable conditions for feasibility. To manage trade-offs, we cast DFL as a multi-objective optimization problem (Miettinen, 1999) and target policies that are Pareto-efficient with respect to the two fairness criteria while achieving high value function. To compute Pareto solutions, we adopt a lexicographic weighted Tchebyshev scalarization strategy (Steuer and Choo, 1983), which can recover Pareto solutions without requiring convexity (see Figure 7 in Appendix). We then develop an estimator that embeds fairness objectives directly into the policy optimization procedure, establish regret bounds that provide rigorous performance guarantees under double fairness constraints.
The proposed DFL framework is flexible and accommodates a broad range of fairness notions from the literature, making it applicable across domains where decisions have heterogeneous downstream impacts. Moreover, DFL offers a practical framework for internal auditing of algorithmic impacts on consumers, helping organizations surface fairness concerns before they escalate into legal or reputational risks. For instance, see the recent action taken by the Federal Trade Commission.
Related Literature. Policy learning seeks to maximize expected reward by learning an optimal policy that selects the best action for each situation. It has broad applications in precision medicine (Murphy, 2003; Robins, 2004; Qian and Murphy, 2011; Zhao et al., 2012; Wallace and Moodie, 2015; Wang et al., 2018; Shi et al., 2018; Qi et al., 2020; Cui and Tchetgen Tchetgen, 2021; Bian et al., 2023), economics (Manski, 2004; Kitagawa and Tetenov, 2018; Athey and Wager, 2021), and business strategy (Den Boer, 2015; Javanmard and Nazerzadeh, 2019; Bian et al., 2024). This literature largely focuses on value maximization, whereas our work augments policy learning with two complementary fairness targets—action fairness (fairness of decisions) and outcome fairness (fairness of the induced consequences)—and studies how to optimize value under both.
Work on fairness-aware policy learning is comparatively scarce and typically enforces a single fairness notion. Nabi et al. (2019) formulate fair policy learning via causal mediation analysis, specifying impermissible causal pathways from the sensitive attribute and optimizing policies under the resulting causal constraints. Tan et al. (2022), Fang et al. (2023), and Cui and Han (2025) develop quantile-optimal decision rules designed to protect disadvantaged groups by controlling tail outcomes. Viviano and Bradic (2023) adopt a “first do no harm” principle by characterizing a welfare Pareto frontier and selecting the fairest allocation within that welfare-efficient set, with computation based on mixed-integer optimization. Wang et al. (2025) study counterfactual fairness in sequential decision making, proposing approaches tailored to reinforcement learning settings. In contrast to these works, we explicitly distinguish action fairness from outcome fairness and address the double fairness problem with different theoretical tools. Figure 1 compares our method to existing fairness-aware policy learning approaches. We group prior single-fairness methods into two categories: action single fair (ASF) methods, which enforce only action fairness, and outcome single fair (OSF) methods, which enforce only outcome fairness. Figure 1 shows that our proposed DFL framework achieves consistently strong performance across all three objectives (action fairness, outcome fairness, and value) with modest compromises in each. By contrast, competing approaches typically improve one dimension at the expense of substantial degradation in the others.
Organization. In Section 2, we provide a brief overview of policy learning and introduce fairness notations for both action fairness and outcome fairness. Section 3 first presents the necessary and sufficient conditions under which a outcome fairness or double fairness policy can be achieved. The task is then formulated within a novel multi-objective optimization framework, and the corresponding estimation procedure is described. Section 4 establishes the theoretical properties of the proposed estimator. Finally, in Sections 5 and 6, we demonstrate the proposed approach using both synthetic and real datasets.
Figure 1: The performance of the three framework types is illustrated using a radar chart, where a higher score along each axis represents better performance in that specific aspect. Here, AF and OF refer to action fairness and outcome fairness, respectively. DFL refers to our proposed double fairness learning method. ASF represents the action single fair policy framework, and OSF denotes the value single fair policy framework.
2 Preliminary
In line with many real-world applications, we consider a binary sensitive attribute and a binary action. Specifically, let $S\in\{0,1\}$ denote the sensitive attribute, $X\in\mathbb{R}^{d}$ a vector of non-sensitive covariates, $A\in\{0,1\}$ the action, and $R$ the resulting outcomes. The data-generating process (DGP) is summarized in Figure 2.
Throughout, we allow for a general two-outcome formulation : $R\equiv(R^{(1)},R^{(2)})^{\top}\in\mathbb{R}^{2}$, where $R^{(1)}$ is the primary outcome that the decision maker seeks to maximize, while $R^{(2)}$ is an outcome along which fairness or equity considerations are assessed. This formulation accommodates a range of applications. In the CASdatasets (Section 6.1), $R^{(1)}$ represents company revenue, while $R^{(2)}$ denotes customer welfare. We also accommodate the degenerate case where $R^{(1)}=R^{(2)}$: for instance, in the government job-training data (Section 6.2), both components correspond to subsequent entrepreneurial activity. In dynamic pricing, $R^{(1)}$ typically represents net profit, whereas $R^{(2)}$ might denote demand or consumer surplus—metrics for which fairness has been extensively studied (see, e.g., Cohen et al., 2022; Bian et al., 2026). Furthermore, in healthcare, $R^{(1)}$ may represent clinical outcomes, while $R^{(2)}$ denotes the allocation of medical resources.
We define $r(s,x,a)=\mathbb{E}\left(R^{(1)}\mid S=s,X=x,A=a\right)$ as the conditional expected primary outcome (the reward function), and $f(s,x,a)=\mathbb{E}\left(R^{(2)}\mid S=s,X=x,A=a\right)$ as the conditional expected fairness-related outcome.
SSXXAARR
Figure 2: The directed acyclic graph of the DGP. The arrow in cyan means that the path is intervenable. The dashed line between $S$ and $X$ means that $S$ might influence $X$, but the relationship is uncertain or not definitively established.
2.1 Policy Learning
We first provide a brief overview of the policy learning task. Recall that $S$ denotes the sensitive attribute, $X$ represents the non-sensitive covariates, $A$ is the action, and $R$ is the reward. A policy $\pi$ prescribes a strategy that selects the action based on $X$ and $S$, which is a function mapping from the information into a probability distribution over the action space. The objective in policy learning is to find an optimal policy $\pi^{\mathrm{opt}}$, such that the value (expected reward) is maximized, i.e.,
$\pi^{\mathrm{opt}}\in\operatorname*{arg\,max}_{\pi\in\Pi}\left[V(\pi)\equiv\mathbb{E}^{\pi}\left(R^{(1)}\right)\right],$
where $\mathbb{E}^{\pi}$ denotes the expectation of the reward where the action follows policy $\pi$, and $\Pi$ is a pre-specified policy class. More specifically, the joint probability of $(X,S,A,R)$ associated with $\mathbb{E}^{\pi}$ is $p(s,x)\pi(a|s,x)p(r|s,x,a)$, where $\pi(a|s,x)$ refers to the probability of choosing the action $a$ given the information $(s,x)$.
A widely used method for estimating $\pi^{\mathrm{opt}}$ involves two steps: policy evaluation and value maximization. In the policy evaluation step, an estimator $\widehat{V}(\pi)$ is obtained for the value function using pre-collected historical data $\{S_{i},X_{i},A_{i},R_{i}\}_{1\leq i\leq n}$, where $n$ is the sample size. One can estimate the value function via the following method:
$\widehat{V}(\pi)=\frac{1}{n}\sum_{i=1}^{n}\sum_{a\in\mathcal{A}}\widehat{r}(S_{i},X_{i},a)\pi(a|S_{i},X_{i}),$
where $\widehat{r}(s,x,a)$ is the regression function estimator of the reward function $r(s,x,a)$. Such a regression approach has been widely used in the policy learning literature; see, e.g., Qian and Murphy (2011); Moodie et al. (2012). Then in the value maximization step, the estimated optimal policy is derived by maximizing this estimated value function, within the pre-specified policy class $\Pi$:
$\widehat{\pi}\in\operatorname*{arg\,max}_{\pi\in\Pi}\widehat{V}(\pi).$
2.2 Fairness Notions
Recently, many fairness notions have been proposed under supervised learning setting, with particular attention to equal opportunity (Hardt et al., 2016) and counterfactual fairness (Kusner et al., 2017). Here, we extend these two notions to the policy learning framework and formally define both action fairness and outcome fairness under each concept.
Definition 1 (Equal Opportunity).
A policy $\pi$ is action-equal opportunity fair if for any $a$, $x$, and any $s\neq s^{\prime}$,
$\pi(a|s,x)=\pi(a|s^{\prime},x).$
A policy $\pi$ is outcome-equal opportunity fair if for any $a$, $x$, and any $s\neq s^{\prime}$,
$\mathbb{E}^{\pi}\left(R^{(2)}\mid s,x\right)=\mathbb{E}^{\pi}\left(R^{(2)}\mid s^{\prime},x\right).$
Equal opportunity requires that individuals with the same observed information $X$ be treated equally across different sensitive groups. This criterion is intuitively appealing and has been widely used in practice (Hardt et al., 2016). However, because it conditions only on observed covariates, it may fail to account for historical or unobserved sources of bias (Chen et al., 2023). To address this, Kusner et al. (2017) proposed counterfactual fairness, which adopts a causal perspective by asking whether the decision (or the induced outcome) would have remained unchanged had the individual belonged to a different sensitive group.
Definition 2 (Counterfactual Fairness).
For an individual with observed attributes $(s,x)$, let $X_{s,x}(s^{\prime})$ denote the counterfactual covariates had the individual’s sensitive attribute been $s^{\prime}$ instead of $s$. A policy $\pi$ is action-counterfactually fair if
$\pi(a|s,X_{s,x}(s))=\pi(a|s^{\prime},X_{s,x}(s^{\prime}))\mbox{ for any }s^{\prime}\neq s,\mbox{ almost surely.}$
and attribute $x$. A policy $\pi$ is outcome-counterfactually fair if for any $s\neq s^{\prime}$,
$\mathbb{E}^{\pi}\left(R^{(2)}|s,X_{s,x}(s)\right)=\mathbb{E}^{\pi}\left(R^{(2)}|s^{\prime},X_{s,x}(s^{\prime})\right)\mbox{ almost surely.}$
Counterfactual fairness leverages counterfactual covariates to account for hidden biases that may not be captured by observed data. Action-counterfactual fairness implies the probability of choosing an action remains invariant under a counterfactual change to the sensitive attribute; and outcome-counterfactual fairness implies that the expected benefit to the individual is independent of their sensitive attribute in a counterfactual world.
To conclude this section, we present the following examples to highlight the distinction between equal opportunity and counterfactual fairness.
Example 1 (Action Fairness).
Equal opportunity. Consider an educational admissions setting where gender is the sensitive attribute $S$ and test score is the observed covariate $X$. An admission decision is action–equal opportunity fair if applicants with the same test score have the same probability of admission, regardless of gender.
Counterfactual fairness. Suppose a female applicant, Nancy, has observed score $X_{s,x}(s)=59$. Now consider a counterfactual world in which Nancy were male; in that world, downstream effects of gender may change her test score to $X_{s,x}(s^{\prime})=50$. An admission decision is action-counterfactually fair if Nancy’s probability of admission in the real world (female, score 59) equals her probability of admission in this counterfactual world (male, score 50).
Example 2 (Outcome fairness (continued)).
Equal opportunity. Suppose the value of an admission policy is defined as the probability of graduating with honors. The policy is outcome–equal opportunity fair if applicants with the same test score have the same probability of graduating with honors, regardless of gender.
Counterfactual fairness. The policy is outcome-counterfactually fair if Nancy’s real-world probability of graduating with honors (female, $X_{s,x}(s)=59$) equals her counterfactual probability (male, $X_{s,x}(s^{\prime})=50$).
2.3 Research Goal
In fairness-aware policy learning, the decision maker’s objectives may be to (a) identify the policy with the highest value function among action fairness policies; (b) find the policy with the highest value function among outcome fairness policies; and (c) find the policy that satisfies both action and outcome fairness criteria while maximizing the value function.
Objective (a) is straightforward, as the policy maker can directly intervene in the action. For example, when considering equal opportunity fairness, one could simply using existing algorithm but impose restriction on the policy class $\{\pi\in\Pi:\pi(a|x)\}$, thereby excluding the sensitive variable for decision making. However, objectives (b) and (c) are more challenging: since direct intervention in the RR(S,X,A)(S,X,A) path is inaccessible, one has to indirectly intervene in AA(S,X)(S,X) to control outcome fairness. Subsequently, several natural questions arise: (i) Does there exist a policy that ensures fairness in its corresponding outcome? And more ambitiously, can a policy achieve double fairness? (ii) If no such fair policy exists, yet the decision maker still aims to mitigate both outcome and action unfairness, how should this be addressed?
In the following section, we explore these questions in depth. Specifically, we address Question (i) in Section 3.1 by establishing necessary and sufficient conditions for the existence of fairness policies. We examine Question (ii) in Section 3.2, where we adopt a multi-objective optimization framework to mitigate both value and action unfairness.
3 Double Fairness Learning
3.1 Existence of Outcome Fairness Policy
Throughout this section, we focus on the equal opportunity fairness, and the discussion of counterfactual fairness is deferred to Section 8.2 of the Appendix. We first present our main results on the existence of outcome fairness and double fairness policies. To begin with, we make the following two assumptions.
Assumption 3.1.
Given a subject with covariates $x$, there exists a treatment option $a^{*}(x)$ that dominates the other treatment with respect to the value function, regardless of the sensitive group, i.e., $\min_{s}f(s,x,a^{*}(x))\geq\max_{s}f(s,x,1-a^{*}(x))$.
Assumption 3.2.
(i) The effects of the sensitive variable have opposite signs across treatments, i.e., $f(1,x,1)-f(0,x,1)][f(1,x,0)-f(0,x,0)]\leq 0$, for all $x$.
(ii) The conditional treatment effects have opposite signs across sensitive variables, i.e., $f(1,x,1)-f(1,x,0)][f(0,x,1)-f(0,x,0)]\leq 0$, for all $x$.
(iii) For all $x$, there exists a pair $(s^{*}(x),a^{*}(x))$ such that $f(s^{*}(x),x,a^{*}(x))\geq f(s^{*}(x),x,1-a^{*}(x))$, and $f(s^{*}(x),x,a^{*}(x))\geq f(1-s^{*}(x),x,a^{*}(x))$.
Both Assumptions 3.1 and 3.2 require that the sensitive variable has only a moderate impact on the resulting value function. Specifically, Assumption 3.1 implies that the conditional treatment effects have the same sign across both sensitive attributes for a given $x$. As changes in the sensitive variable within the reward function do not alter the sign of the treatment effects, it suggests that the effects of the sensitive variable are relatively mild compared to the treatment effects. As demonstrated in Proposition 1, if the effects of the sensitive variable are relatively limited, a outcome fairness policy exists, which is intuitively reasonable. If Assumption 3.1 does not hold (i.e., Assumption 3.2 (ii) applies), achieving outcome fairness requires additional regularity conditions on the effects of the sensitive variable. Specifically, Assumption 3.2 (i) requires that the effects of the sensitive variable have opposite signs across treatments for a given $x$, meaning that no sensitive group dominates the other under either treatment choice. Furthermore, when Assumption 3.2 (i) is satisfied, a unique double fairness policy exists (refer to Proposition 2). We next summarize the main results.
Proposition 1.
There exists a outcome fairness policy for the given environment if and only if either Assumption 3.1 or Assumption 3.2 holds.
Proposition 2.
A double fairness policy is either almost surely unique or does not exist. Furthermore, a double fairness policy exists if and only if Assumption 3.2 (i) holds.
Propositions 1 and 2 establish the sufficient and necessary conditions for the existence of the outcome fairness policy and the double fairness policy, respectively. A similar result holds for counterfactual fairness; see Section 8.3 of the Supplemental Materials, which also discusses extensions of these results to multi-action settings beyond the binary case.
Since a outcome fairness or a double fairness policy may not always exist when performing the policy learning, one could instead seek a policy that is as fair as possible in both value function and action. In the next section, we present a multi-objective policy learning approach to achieve this goal. Notably, if a doubly fairness policy exists, our proposed method below can automatically identify it.
3.2 Multi-objective Policy Learning
Since achieving exact action and outcome fairness is generally challenging, we first introduce fairness metrics to quantify the degree of unfairness. Specifically, let $\Delta_{1}(\pi)$ and $\Delta_{2}(\pi)$ denote generic metrics for action fairness and outcome fairness, respectively, where smaller values indicate less unfairness. Below, we present several examples of fairness metrics using equal opportunity and counterfactual fairness. For brevity, in what follows, we write $\pi(s,x)$ to represent $\pi(1|s,x)$.
Example 3 (Equal Opportunity).
For equal opportunity fairness, a fairness metric for either action or value can be defined as following:
$\Delta_{1}(\pi)=\mathbb{E}\left[\pi(1,X)-\pi(0,X)\right]^{2}\mbox{ and }\Delta_{2}(\pi)=\mathbb{E}\left[f^{\pi}(1,X)-f^{\pi}(0,X)\right]^{2},$
where $f^{\pi}(s,x)=\mathbb{E}^{\pi}\left(R^{(2)}\mid s,x\right)$. They can be estimated using
$\widehat{\Delta}_{1}(\pi)=\frac{1}{n}\sum_{i=1}^{n}\left[\pi(1,X_{i})-\pi(0,X_{i})\right]^{2}\mbox{ and }\widehat{\Delta}_{2}(\pi)=\frac{1}{n}\sum_{i=1}^{n}\left[\widehat{f}^{\pi}(1,X_{i})-\widehat{f}^{\pi}(0,X_{i})\right]^{2},$
where $\widehat{f}^{\pi}(s,x)$ is some estimator for $f^{\pi}(s,x)$.
Example 4 (Counterfactual Fairness).
A counterfactual fairness metric for action or value can be defined as:
$\Delta_{1}(\pi)=\mathbb{E}\left[\pi(S,X)-\pi(S^{\prime},X^{\prime})\right]^{2}\mbox{ and }\Delta_{2}(\pi)=\mathbb{E}\left[f^{\pi}(S,X)-f^{\pi}(S^{\prime},X^{\prime})\right]^{2},$
where $S^{\prime}=1-S$, and $X^{\prime}\equiv X_{S,X}(S^{\prime})$. They can be estimated by
$\widehat{\Delta}_{1}(\pi)=\frac{1}{n}\sum_{i=1}^{n}\left[\pi(S_{i},X_{i})-\pi(S_{i}^{\prime},\widehat{X}_{i}^{\prime})\right]^{2},\mbox{ and }\widehat{\Delta}_{2}(\pi)=\frac{1}{n}\sum_{i=1}^{n}\left[\widehat{f}^{\pi}(S_{i},X_{i})-\widehat{f}^{\pi}(S_{i}^{\prime},\widehat{X}_{i}^{\prime})\right]^{2}.$
Note that here the counterfactual covariates $X_{i}^{\prime}$ need to be estimated by $\widehat{X}_{i}^{\prime}$. The entire estimation procedure will be discussed in Section 4.1.
Making optimal decisions in the presence of trade-offs among multiple objectives is inherently challenging, as it is generally impossible to optimize all objectives simultaneously. Therefore, it is crucial to understand how to effectively balance multiple objectives and make optimal trade-offs among them. The concept of Pareto set (see, e.g., Miettinen, 1999) offers a valuable framework for this purpose, capturing the idea that improvements in one objective cannot be achieved without compromising at least one other. By leveraging the idea of Pareto set, we can navigate complex decision-making scenarios and strive for policies that optimize overall benefits across diverse objectives.
We now incorporate the two previously defined fairness metrics to formulate the problem as a multi-objective policy learning task. Specifically, our proposal involves identifying the optimal Pareto fairness policy within the Pareto fairness set $\Pi_{p}$ that maximizes the value function:
$\pi^{*}\in\operatorname*{arg\,max}_{\pi\in\Pi_{p}}V(\pi).$ (1)
Here, $\Pi_{p}$ denotes the set of Pareto fairness policies–those for which no other policy achieves strictly better action fairness and outcome fairness simultaneously. We formally define the Pareto fairness set for our problem below.
Definition 3 (Pareto Fairness Policies and Pareto Fairness Set).
Let $\Delta_{1}(\pi)$ and $\Delta_{2}(\pi)$ denote generic fairness metrics for measuring the degree of action and outcome fairness, respectively. A policy $\pi_{p}$ is said to be a Pareto fairness policy if there does not exist another policy $\pi^{\prime}$ such that $\Delta_{1}(\pi^{\prime})\leq\Delta_{1}(\pi_{p})$ and $\Delta_{2}(\pi^{\prime})\leq\Delta_{2}(\pi_{p})$, with at least one of these two inequalities being strict. The Pareto fairness set, denoted by $\Pi_{p}$, is the collection of all such Pareto policies.
Remark 1.
Recall from Proposition 2 that if a double fairness policy $\bar{\pi}$ exists (i.e., $\Delta_{1}(\bar{\pi})=\Delta_{2}(\bar{\pi})=0$), then it is almost surely unique. Since the fairness metrics are non-negative, no other policy $\pi^{\prime}$ satisfies $\Delta_{1}(\pi^{\prime})<\Delta_{1}(\bar{\pi})$ or $\Delta_{2}(\pi^{\prime})<\Delta_{2}(\bar{\pi})$. Thus, by Definition 3, $\Pi_{p}$ must be a singleton containing only $\bar{\pi}$. Consequently, if a double fairness policy exists, our proposed method in Equation (1) will automatically identify it.
Next we demonstrate how to estimate $\pi^{*}$ as defined in Equation (1). Naturally, the estimation procedure consists of first approximating the Pareto set $\Pi_{p}$ by an estimated Pareto set $\widehat{\Pi}_{p}$, followed by choosing the policy within $\widehat{\Pi}_{p}$ that maximizes the estimated value function. Although Definition 3 offers an intuitive understanding of the Pareto fairness set, directly estimating the fairness Pareto set according to this definition can be challenging. To better characterize the Pareto fairness set, we present the following useful proposition.
Proposition 3 (Lexicographic Weighted Tchebychef Method (Steuer and Choo, 1983)).
The Pareto fairness set can be computed via
$\Pi_{p}=\bigcup_{\alpha\in(0,1)}\big\{\pi:\operatorname*{arg\,inf}_{\pi\in\Lambda_{\alpha}^{*}}\Delta(\pi)\big\},\mbox{ where }\Delta(\pi)=\Delta_{1}(\pi)+\Delta_{2}(\pi),$
$\Lambda_{\alpha}^{*}=\big\{\pi:\operatorname*{arg\,inf}_{\pi\in\Pi}M_{\alpha}(\pi)\big\},\mbox{ and }M_{\alpha}(\pi)=\max\left\{\alpha\Delta_{1}(\pi),(1-\alpha)\Delta_{2}(\pi)\right\},$
where $\Pi$ is a pre-specified policy class.
Remark 2.
Viviano and Bradic (2023) adopted the linear scalarization approach (specific formulation discussed in Section 5) to characterize the Pareto set. However, this method cannot identify Pareto fairness policies in the non-convex regions of the Pareto fairness front (as illustrated in Figure 7 in Section 8.1 of the Supplemental Materials). By contrast, the lexicographic weighted Tchebycheff approach can identify the full set of Pareto fairness policies without requiring convexity. We also include a simple intuitive example in Section 8.1 to demonstrate that the linear scalarization approach fails, while the weighted Tchebycheff method works.
Now it remains to estimate the Pareto fairness set via Proposition 3. The estimation procedure consists of the following steps. First, the fairness metrics $\Delta_{1}(\pi)$ and $\Delta_{2}(\pi)$ are estimated by $\widehat{\Delta}_{1}(\pi)$ and $\widehat{\Delta}_{2}(\pi)$, respectively. Then for any fixed $\alpha\in(0,1)$, the set $\Lambda_{\alpha}^{*}$ is estimated by
$\widehat{\Lambda}_{\alpha}^{*}=\left\{\pi\in\Pi:\widehat{M}_{\alpha}(\pi)\leq\widehat{M}_{\alpha}^{*}+\kappa\right\},$ (2)
 $\mbox{ where }\widehat{M}_{\alpha}(\pi)=\max\left\{\alpha\widehat{\Delta}_{1}(\pi),(1-\alpha)\widehat{\Delta}_{2}(\pi)\right\},\mbox{ and }\widehat{M}_{\alpha}^{*}=\inf_{\pi\in\Pi}\widehat{M}_{\alpha}(\pi).$
Here, $\kappa$ is a slack parameter, chosen to be a sufficiently small constant that ensures $\Lambda_{\alpha}^{*}\subseteq\widehat{\Lambda}_{\alpha}^{*}$ with high probability. The Pareto fairness set then can be estimated by
$\widetilde{\Pi}_{p}=\bigcup_{\alpha\in(0,1)}\left\{\pi\in\widehat{\Lambda}_{\alpha}^{*}:\widehat{\Delta}(\pi)\leq\widehat{\Delta}_{\alpha}^{*}+\kappa\right\},$
where $\widehat{\Delta}(\pi)=\widehat{\Delta}_{1}(\pi)+\widehat{\Delta}_{2}(\pi)$, and $\widehat{\Delta}_{\alpha}^{*}=\inf_{\pi\in\widehat{\Lambda}_{\alpha}^{*}}\widehat{\Delta}(\pi)$. Similarly, we introduce a slack parameter to ensure that the approximated Pareto fairness set $\widetilde{\Pi}_{p}$ covers the true set $\Pi_{p}$ with high probability. Note that we can use the same slack parameter $\kappa$ as before, since the convergence rates of $\widehat{M}_{\alpha}(\pi)$ and $\widehat{\Delta}(\pi)$ are of the same order (as shown in Lemma 1 in the supplemental material).
Since deriving the approximated set $\widetilde{\Pi}_{p}$ requires searching over all $\alpha\in(0,1)$, we discretize the interval $(0,1)$ by defining $\Omega=\{\alpha_{1},\alpha_{2},\dots,\alpha_{K}\}$, where $K$ denotes a pre-determined discretization level. Moreover, we require that for any $\alpha\in(0,1)$, there exists $\alpha_{k}\in\Omega$ such that $|\alpha-\alpha_{k}|\leq K^{-1}$. We show later in Section 4 that it suffices to set $K=O(\sqrt{n/\log n})$ to guarantee desirable theoretical properties. Then we have
$\widehat{\Pi}_{p}=\bigcup_{k=1}^{K}\left\{\pi\in\widehat{\Lambda}_{\alpha_{k}}^{*}:\widehat{\Delta}(\pi)\leq\widehat{\Delta}_{\alpha_{k}}^{*}+\kappa\right\}.$
Finally, the estimated policy is
$\widehat{\pi}\in\operatorname*{arg\,max}_{\pi\in\widehat{\Pi}_{p}}\widehat{V}(\pi).$ (3)
We summarize the estimation procedure in Algorithm 1.
Algorithm 1 Pseudocodes for Estimating $\widehat{\pi}$.
1:function ($\{S_{i},X_{i},A_{i},R_{i}\}_{1\leq i\leq n},\;\Pi,\kappa,K$)
2:  Calculate $\widehat{\Delta}_{1}(\pi)$, $\widehat{\Delta}_{2}(\pi)$, and $\widehat{V}(\pi)$, for $\pi\in\Pi$.
3:  for $k=1,2\dots,K$ do
4:   Compute $\widehat{M}_{\alpha_{k}}^{*}=\inf_{\pi\in\Pi}\widehat{M}_{\alpha_{k}}(\pi)=\inf_{\pi\in\Pi}\max\{\alpha_{k}\widehat{\Delta}_{1}(\pi),(1-\alpha_{k})\widehat{\Delta}_{2}(\pi)\}$.
5:   Compute $\widehat{\Delta}_{\alpha_{k}}^{*}=\inf_{\pi\in\Pi}\widehat{\Delta}(\pi)$, s.t. $\widehat{M}_{\alpha_{k}}(\pi)\leq\widehat{M}_{\alpha_{k}}^{*}+\kappa$.
6:   Compute $\widehat{V}_{k}^{*}=\sup_{\pi\in\Pi}\widehat{V}(\pi)$, s.t. $\widehat{\Delta}(\pi)\leq\widehat{\Delta}_{\alpha_{k}}^{*}(\pi)+\kappa,\mbox{ and }\widehat{M}_{\alpha_{k}}(\pi)\leq\widehat{M}_{\alpha_{k}}^{*}+\kappa$.  Denote the solution as $\widehat{\pi}_{k}^{*}$.
7:  Compute $k^{*}=\operatorname*{arg\,max}_{1\leq k\leq K}\widehat{V}_{k}^{*}$ .
8:  return $\widehat{\pi}=\widehat{\pi}_{k^{*}}^{*}$.
4 Theoretical Results
In this section, we first establish the theoretical properties of our method under general fairness notion and metric. Theorem 1 provides the regret bound for the estimated policy in terms of the Hausdorff distance between the estimated and true Pareto sets; while Theorem 2 further bounds this distance. Both results hold for any fairness metric within our framework. We then specialize these results to equal opportunity and counterfactual fairness, deriving the corresponding regret bounds in Theorems 3 and 4.
We begin by defining the regret function. The regret of $\widehat{\pi}$ quantifies the performance gap between $\pi^{*}$ and $\widehat{\pi}$, and is defined as
[…]
Assumption 4.1.
The functions $r(s,x,a)$, $\widehat{r}(s,x,a)$, $f(s,x,a)$, and $\widehat{f}(s,x,a)$ are uniformly bounded in absolute value by $R_{\max}$.
Theorem 1 (Regret Bound).
Suppose Assumption 4.1 holds, and the slack parameter $\kappa$ satisfies
$\kappa\geq C\max\Big[\sup_{\pi\in\Pi}\big|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\big|+\sup_{\pi\in\Pi}\big|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\big|,\;K^{-1}\Big],$
[…]
$-H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})\;\lesssim\;\mu(\widehat{\pi})\;\lesssim\;\min\Big[\sup_{\pi\in\Pi_{p}}|\eta(\pi)|+H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p}),\;\sup_{\pi\in\Pi}|\eta(\pi)|\Big],$
almost surely. Here, $\eta(\pi)\equiv\widehat{V}(\pi)-V(\pi)$ denotes the policy evaluation error, and $H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})$ is the Hausdorff distance between $\Pi_{p}$ and $\widehat{\Pi}_{p}$ under the total variation metric [^2] Here we use the total variation distance $d_{\mathrm{TV}}$, so that $H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})=\max\Big\{\sup_{\pi\in\Pi_{p}}\inf_{\hat{\pi}\in\widehat{\Pi}_{p}}d_{\mathrm{TV}}(\pi,\hat{\pi}),\;\sup_{\hat{\pi}\in\widehat{\Pi}_{p}}\inf_{\pi\in\Pi_{p}}d_{\mathrm{TV}}(\hat{\pi},\pi)\Big\},$ $d_{\mathrm{TV}}(\pi,\hat{\pi}):=\sup_{s,x}d_{\mathrm{TV}}\big(\pi(\cdot\mid s,x),\hat{\pi}(\cdot\mid s,x)\big)=\sup_{s,x}\frac{1}{2}\sum_{a\in\mathcal{A}}\big|\pi(a\mid s,x)-\hat{\pi}(a\mid s,x)\big|.$ .
Theorem 1 shows the regret consists of two parts: (i) the uniform policy evaluation error over $\Pi_{p}$ and $\Pi$, and (ii) the estimation error of the Pareto fairness set, measured by the Hausdorff distance $H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})$. For (i), if the nuisance functions are estimated with sufficient accuracy and $\Pi$ is a VC subgraph class, then the uniform error admits a suitable bound; see, e.g., Kitagawa and Tetenov (2018); Athey and Wager (2021). We now analyze component (ii), starting with the following assumption.
Assumption 4.2 (Well-separated Pareto fairness Set).
Let $\Pi_{\delta}\equiv\left\{\pi\in\Pi:d_{\mathrm{TV}}(\pi,\Pi_{p})>\delta\right\}$ denote the set of policies whose total variation distance from the true Pareto set is larger than $\delta$, where $\delta$ is a positive constant. Assume that for all $\alpha\in(0,1)$, at least one of the following conditions holds:
$(i)\inf_{\pi\in\Pi_{\delta}}\left|\Delta(\pi)-\Delta_{\alpha}^{*}\right|>c\delta;\quad(ii)\inf_{\pi\in\Pi_{\delta}}\left|M_{\alpha}(\pi)-M_{\alpha}^{*}\right|>c\delta,$
for some universal constant $c>0$.
Assumption 4.2 imposes a separation condition: policies that are at least $\delta$ away from $\Pi_{p}$ in total variation must incur an objective gap of order $\delta$ under either the criterion $\Delta(\cdot)$ or the weighted Tchebycheff criterion $M_{\alpha}(\cdot)$. Such well-separated solution conditions are standard in establishing consistency of $M$- and $Z$-estimators; see, e.g., Van der Vaart (2000).
Theorem 2.
Suppose that the conditions in Theorem 1 and Assumption 4.2 hold, then
$H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})\lesssim\kappa,$
almost surely, where $\kappa$ is the slack parameter in Algorithm 1.
Theorem 2 shows that, up to a constant factor, the Hausdorff distance between the estimated and true Pareto fairness sets is bounded by $\kappa$. Recall that $\kappa$ is determined by two terms: the estimation error in the fairness metrics and the discretization error $K^{-1}$. If $K^{-1}$ is chosen to be asymptotically negligible relative to the estimation error, then the resulting bound is driven by the estimation error.
Building on the general results above, the following subsection presents an in-depth analysis of equal opportunity and counterfactual fairness.
4.1 Equal Opportunity and Counterfactual Fairness Applications
In this subsection, we demonstrate how our theoretical results apply to the notions of equal opportunity and counterfactual fairness. We begin by imposing some technical conditions.
Assumption 4.3 (Finite VC Subgraph Dimension).
The policy class $\Pi$ has finite VC-subgraph dimension; that is, $\mbox{VC}(\Pi)$ is bounded.
Assumption 4.3 is commonly imposed in policy learning literature, see, e.g., Kitagawa and Tetenov (2018); Athey and Wager (2021). Common examples of classes with finite VC-subgraph dimension include linear decision rules, decision trees, and sigmoid (logistic) score classes.
Assumption 4.4 (Nuisance’s Regularity Conditions).
There exists $\gamma_{1},\;\gamma_{2}\in(0,1]$ such that the estimated reward function satisfies
$\max_{a,s}\;\frac{1}{n}\sum_{i=1}^{n}\left[\widehat{r}(s,X_{i},a)-r(s,X_{i},a)\right]^{2}=O_{p}(n^{-\gamma_{1}}),$
and $\max_{a,s}\;\frac{1}{n}\sum_{i=1}^{n}\left[\widehat{f}(s,X_{i},a)-f(s,X_{i},a)\right]^{2}=O_{p}(n^{-\gamma_{2}}).$
Assumption 4.4 is needed because the outcome fairness metric estimator $\widehat{\Delta}_{2}(\pi)$ and the value estimator $\widehat{V}(\pi)$ are constructed from $\widehat{r}$ and $\widehat{r}$, respectively. The exponents $\gamma_{1}$ and $\gamma_{2}$ depend on the methods used to estimate $r$ and $f$. If $r$ and $f$ are estimated using correctly specified parametric models, one typically has $\gamma_{1}=\gamma_{2}=1$. Alternatively, flexible nonparametric methods such as sieve estimators (Huang, 1998) can be used. Under standard smoothness and complexity conditions, sieve estimators can achieve the optimal mean-squared rate $O_{p}(n^{-\frac{2\beta}{2\beta+d}})$, where $\beta$ is the smoothness parameter of the reward function, and $d$ is the dimension of $X$. Sieve methods are widely used in the policy evaluation and learning literature; see, e.g., Shi et al. (2022); Chen and Qi (2022); Bian et al. (2025).
For the counterfactual fairness notion, we additionally impose the following Assumption 4.5 to ensure that the counterfactual covariates $X_{s,x}(s^{\prime})$ are identifiable from the observed data. Similar assumptions are used in Chen et al. (2023) and Wang et al. (2025) to study counterfactual fairness.
Assumption 4.5.
For each $j\in\{1,\ldots,d\}$, the covariate satisfies $X_{j}=\theta_{j}(S)+\varepsilon_{j}$, where $\theta_{j}(\cdot)$ is the regression function for the $j$th covariate, and the error term $\varepsilon_{j}$ satisfies $\mathbb{E}(\varepsilon_{j}|S)=0$.
Under Assumption 4.5, for an individual with observed $(S,X)=(s,x)$, the counterfactual covariates under $S=s^{\prime}$ are
$X_{s,x}(s^{\prime})=\theta(s^{\prime})+x-\theta(s),\mbox{ with }\theta(\cdot)=\big(\theta_{1}(\cdot),\ldots,\theta_{d}(\cdot)\big)^{\top}.$
We next impose a Lipschitz-type condition on the estimated reward function and the policy class. This ensures that small errors in estimating the counterfactual covariates lead to correspondingly small perturbations in the quantities evaluated at $(s^{\prime},x^{\prime})$.
Assumption 4.6 (Lipschitz Condition).
There exists a constant $m>0$ such that
$(i)\;\;|\widehat{f}(s,x,a)-\widehat{f}(s,x+t,a)|\leq m\lVert t\rVert_{2},\mbox{ for any }s,a,\mbox{ and }x;$
$(ii)\sup_{\begin{subarray}{c}\pi\in\Pi\\ t:\,\lVert X-t\rVert_{2}\leq\delta\end{subarray}}\mathbb{E}\left[\pi(S,X)-\pi(S,t)\right]^{2}\leq m\delta.$
Assumption 4.6(i) is standard and has also been imposed by Wang et al. (2025) for counterfactual fairness policy learning. Assumption 4.6(ii) requires only Lipschitz continuity in expected mean square, which is mild enough to accommodate deterministic policies (for which pointwise Lipschitz continuity typically fails). A closely related condition is used by Shi et al. (2021) when analyzing linear deterministic decision rules; see Assumption A5(ii) therein.
We now present the main regret bounds for equal opportunity and counterfactual fairness.
Theorem 3 (Equal Opportunity).
Suppose Assumptions 4.1, 4.2, 4.3, and 4.4 hold. In addition, set the slack parameter $\kappa$ to $c_{0}\max(n^{-\frac{\gamma_{2}}{2}},\sqrt{\log n/n})$ for some constant $c_{0}>0$, and that $K^{-1}=O(\sqrt{\log n/n})$. Then under the equal opportunity notion of fairness,
$|\mu(\widehat{\pi})|=O_{p}\big(\sqrt{\mbox{VC}(\Pi)/n}+\sqrt{\log n/n}\big)+O_{p}\big(n^{-\frac{\gamma_{1}}{2}}+n^{-\frac{\gamma_{2}}{2}}\big).$
Theorem 4 (Counterfactual).
Suppose the assumptions of Theorem 3 hold. In addition, assume Assumptions 4.5 and 4.6. Then, under the counterfactual notion of fairness, we have
$|\mu(\widehat{\pi})|=O_{p}\big(\sqrt{\mbox{VC}(\Pi)/n}+\sqrt{\log n/n}+\sqrt{d/n}\big)++O_{p}\big(n^{-\frac{\gamma_{1}}{2}}+n^{-\frac{\gamma_{2}}{2}}\big).$
Theorems 3 and 4 establish the regret bound of the estimated policy under the notions of equal opportunity fairness and counterfactual fairness, respectively. Specifically, the terms $O_{p}(\sqrt{\text{VC}(\Pi)/n})$ and $O_{p}(\sqrt{\log n/n})$ arise from the empirical process argument and concentration inequality, whereas the terms $O_{p}(n^{-\frac{\gamma_{1}}{2}})$ and $O_{p}(n^{-\frac{\gamma_{2}}{2}})$ reflects the estimation error in the outcome. Relative to equal opportunity, the counterfactual fairness bound includes an additional $\sqrt{d/n}$ term, stemming from the extra step of estimating counterfactual covariates.
Remark 3.
In Theorems 3 and 4, the slack parameter $\kappa$ involves the term $n^{-\frac{\gamma_{2}}{2}}$. In practice, if a parametric model is used, one can typically choose $\gamma_{2}=1$. For nonparametric regression methods such as the sieve estimator, $\gamma_{2}$ is given by $\frac{2\beta}{2\beta+d}$, as noted earlier. Since $\beta$ is typically unknown, one may infer it from the optimal complexity choice. For sieve estimation, the optimal number of basis functions satisfies $L^{\mathrm{opt}}\;=\;n^{\frac{d}{d+2\beta}}.$ Thus, one can first select $L^{\text{opt}}$ (e.g., via cross-validation), and then back out $\beta$ and hence $\gamma_{2}$ accordingly. The same logic extends to other nonparametric methods (e.g., kernel or local polynomial estimators), whose optimal bandwidth choices also depend on $\beta$; see, for example, Tsybakov (2009).
5 Numerical Studies
In this section, we demonstrate the effectiveness of our proposed framework through extensive simulation studies. We begin by outlining the data generating process in detail.
Data generating procedure. We consider the following data-generating process, which is similar to that in Section F.2.1 of Wang et al. (2025). The sensitive variable is given by $S_{i}\sim\mbox{Bernoulli}(0.35)$, and we consider the case in which the covariates are two-dimensional and are generated by $X_{1i}=0.45S_{i}+\varepsilon_{1i}$, and $X_{2i}=0.85S_{i}+\varepsilon_{2i}$, where $\{\varepsilon_{1i}\}_{i}$ and $\{\varepsilon_{2i}\}_{i}$ are independent error terms drawn from a standard Gaussian distribution. The action is sampled using $A_{i}\sim\mbox{Bernoulli}(\operatorname*{expit}(-0.5+S_{i}+X_{1i}-X_{2i}))$, where $\operatorname*{expit}(t)=\exp(t)/(1+\exp(t))$. Finally, we consider the case where $R=R^{(1)}=R^{(2)}$, such that $R$ is one-dimensional. In this setting, the reward is generated according to: $R_{i}=1.5S_{i}+0.9X_{1i}+0.8X_{2i}+A_{i}(1-0.8S_{i}-0.5X_{1i}-0.5X_{2i})+\epsilon_{i},$ where $\{\epsilon_{i}\}_{i}$ are independent error terms drawn from a standard Gaussian distribution. Under this setting, the global optimal decision rule is a linear decision rule $\mathds{1}(1-0.8S_{i}-0.5X_{1i}-0.5X_{2i}>0)$.
Competing methods. We compare our proposed method with the standard policy learning approach, and the approach for finding Pareto policies in Viviano and Bradic (2023). Specifically, the standard policy learning approach estimates the optimal policy as
$\operatorname*{arg\,max}_{\pi\in\Pi}\;\frac{1}{n}\sum_{i=1}^{n}\left[\widehat{r}(S_{i},X_{i},1)-\widehat{r}(S_{i},X_{i},0)\right]\pi(S_{i},X_{i}).$
As for the approach outlined in Viviano and Bradic (2023), it focuses solely on single fairness policy learning tasks, either outcome fairness or action fairness. To build on this, we first consider the original method based on the action fairness metric (VB1) and the outcome fairness metric (VB2). Moreover, we consider an adaptive version (ADVB), which has been tailored to accommodate the double fairness setting. Specifically, the ADVB method is adapted to our proposed double fairness setting utilizing the linear scalarization method to derive the resulting policy. As in our proposed procedure, it involves two steps: first, the estimated Pareto set $\widehat{\Pi}_{ADVB}$ is obtained using linear scalarization, i.e.,
$\left\{\pi\in\Pi:\alpha_{k}\widehat{\Delta}_{1}(\pi)+(1-\alpha_{k})\widehat{\Delta}_{2}(\pi)\leq\min_{\pi}\left(\alpha_{k}\widehat{\Delta}_{1}(\pi)+(1-\alpha_{k})\widehat{\Delta}_{2}(\pi)\right)+\kappa,\forall\alpha_{k}\in\Omega\right\}.$
Subsequently, the estimated policy is obtained by maximizing the estimated value function within $\widehat{\Pi}_{ADVB}$: $\operatorname*{arg\,max}_{\pi\in\widehat{\Pi}_{ADVB}}\widehat{V}(\pi).$ Throughout this section, we apply a linear decision rule across all methods and evaluate their performance on a test set of 5,000 samples using both equal opportunity fairness and counterfactual fairness metrics. In addition, the value estimator is computed using a correctly specified outcome regression approach for all methods.
Figure 3: Empirical action fairness metric $\Delta_{1}(\pi)$ (lower is better), outcome fairness metric $\Delta_{2}(\pi)$ (lower is better), and value function $V(\pi)$ (higher is better), based on the equal opportunity fairness notion. Results are obtained from a testing set of size 5,000, comparing our proposed DFL method with competing approaches. The results are averaged over 100 replications, each with $K=10$ and a training set of size 200.
In the first part of our experiments, we investigate the effect of the slack parameter $\kappa$. According to Theorems 3 and 4, it is sufficient to set $\kappa=c_{0}\sqrt{\log n/n}$ for some constant $c_{0}$, given that the reward function is estimated using a correctly specified model. In the following, we consider four scenarios corresponding to different values of $c_{0}$: $0.2,0.5,1,$ and $2$, respectively.
The simulation results for both equal opportunity and counterfactual fairness are summarized in Figures 3 and 4, respectively. As evidenced by the results, our proposed method is not sensitive to the choice of $\kappa$. Moreover, across all scenarios, the proposed DFL method demonstrates superior performance. For instance, under the equal opportunity fairness notion, the DFL policy achieves a value function comparable to the unconstrained optimal policy while significantly enhancing fairness in both actions and value. Notably, our method achieves near-zero action fairness metrics for $c_{0}=0.2$; and its outcome fairness metrics are comparable to those achieved by the VB2 method, which focuses solely on minimizing outcome fairness. This underscores the DFL method’s ability to effectively balance all three objectives—without compromising overall welfare or revenue.
Figure 4: Empirical action fairness metric $\Delta_{1}(\pi)$ (lower is better), outcome fairness metric $\Delta_{2}(\pi)$ (lower is better), and value function $V(\pi)$ (higher is better) based on the counterfactual fairness notion, obtained from a testing set of size 5,000, comparing our proposed DFL method with competing approaches. The results are averaged over 100 replications, each with $K=10$ and a training set of size 200.
Figure 5: Performance, based on numerical studies, is visualized using a radar chart. A higher score along a particular axis indicates better performance in that aspect. DFL refers to our proposed double fairness learning method. ASF represents the action single fair policy framework (VB1), and OSF denotes the outcome single fair policy framework (VB2).
In contrast, the ADVB method results in a lower value function compared to DFL and exhibits higher action and outcome fairness metrics. Similarly, while the VB1 and VB2 methods prioritize single fairness criteria (action fairness for VB1 and outcome fairness for VB2), they perform poorly in terms of value function (both VB1 and VB2), outcome fairness (VB1), and action fairness (VB2). Overall, only the proposed DFL method achieves a flexible and effective balance among the three objectives, demonstrating both practical utility and adaptability. Figure 5 summarizes these results using a radar chart.
Figure 6: Empirical action fairness metric $\Delta_{1}(\pi)$ (lower is better), outcome fairness metric $\Delta_{2}(\pi)$ (lower is better), and value function $V(\pi)$ (higher is better), based on two fairness notions, evaluated from a testing set of size 5,000, using our proposed DFL method. The results are averaged over 100 replications, each with $\kappa=0.5\sqrt{\log n/n}$ and a training set of size 200.
In the previous results, the discretization size $K$ was fixed at 10. In the second part of our experiments, we investigate the impact of different $K$ values on performance. For brevity, we only present our proposed DFL method, as the performance of all methods remains largely unaffected by changes in $K$. As shown in Figure 6, varying the value of $K$ leads to negligible changes in performance.
6 Real Data Analysis
6.1 Motor Insurance Liability Claims Data
In this section, we illustrate our proposed method to a Belgian motor third-party liability dataset from the CASdatasets (Dutang and Charpentier, 2019). Similar to Xin and Huang (2024), we use gender as the sensitive variable. In the offline dataset, we classify customers as low-risk or high-risk based on the bonus-malus (BM) scale, following common practice in the insurance literature, where drivers are often categorized as good or bad based on their BM score (Frangos and Vrontos, 2001). The BM scale ranges from 0 to 22, with higher values indicating greater risk. We define customers with BM greater than five as high-risk and the rest as low-risk. The dataset consists of 18,276 observations, with 28% being female and 35% classified as high-risk. The non-sensitive covariates include the number of claims, policyholder age, coverage plan, if the vehicle is part of a fleet, and the coverage duration. Since the dataset does not contain premium information, we generate the pure premium amount using the frequency-severity approach, as outlined by Baumann and Loi (2023). This approach sets pure premiums based on the predicted risk. To better reflect real-world pricing, we further adjust the final premium based on the coverage plan and gender. See Section 8.4 in Supplemental Materials for more details. Accordingly, the benefit or welfare received by the customer is the negative of the reward. Hence, it is essential to consider both action and outcome fairness in this setting.
Table 1: The estimated value function, the estimated action fairness metric, and the estimated outcome fairness metric using our proposed DFL and other competing methods.
Equal Opportunity Counterfactual
[…]
DFL -10.16 0.03 0.18 -10.18 0.02 0.12
Optimal -9.38 0.33 1.14 -9.38 0.38 1.10
VB1 -10.22 0.00 0.40 -10.40 0.00 0.49
VB2 -10.34 0.05 0.06 -10.23 0.04 0.11
ADVB -10.17 0.00 0.48 -10.40 0.00 0.49
We illustrate our method for both equal opportunity fairness, and the counterfactual fairness. Throughout, we set $K=10$ and $\kappa=2\sqrt{\log n/n}$. Table 1 provides a summary of results. Similar to findings from previous numerical studies, our proposed policy yields a estimated value function similar to the estimated optimal policy, while being much fairer than the optimal policy in terms of both action fairness and outcome fairness. This demonstrates that the newly proposed approach effectively balances the three objectives without compromising the overall revenue.
6.2 Entrepreneurship Training Data
In this subsection, we illustrate our proposed method to the entrepreneurship training data for North America undergraduate students from Lyons and Zhang (2017). This study examined whether entrepreneurship programs have varying impacts on certain minority groups, such as females or non-Caucasians. Focusing solely on the outcome fairness criterion, they found that the program’s short-term effect on minorities is minimal. The goal in our study is to identify the policy that maximizes the expected subsequent entrepreneurial activity for the population while keeping the action and value unfairness as low as possible.
In this context, the action represents the decision of whether to accept or reject admission into the program; while the reward is defined as subsequent entrepreneurial activity, specifically whether finalists work in the startup sector during the period immediately following the program. We focus on the sensitive variable gender. The covariates in this dataset include the average interview score, prior entrepreneurial activity, and school rank. The dataset includes a total of 335 subjects, with 53.43% admitted and the remaining not admitted. Additionally, 26.57% of the subjects are female; among those who were admitted, 27.93% are female.
Since the reward is binary, we fit the outcome model using logistic regression. Note that during the training process, the optimal policy for this dataset turns out to be a one-size-fits-all approach, where ideally, everyone would be admitted to the program. This makes the action exactly fair, placing the optimal policy within our Pareto efficiency set. To address this issue, we introduce an additional constraint on the pre-defined policy class, ensuring that the estimated number of admitted individuals does not exceed the number of admitted individuals in the dataset (179 in total). This constraint is practical, as many real-world applications face limitations in resources and budget.
Table 2: The estimated value function, the estimated action fairness metric, and the estimated outcome fairness metric using our proposed DFL and other competing methods.
Equal Opportunity Counterfactual
[…]
DFL 0.22 0.09 0.06 0.22 0.09 0.07
Optimal 0.23 0.68 0.12 0.23 0.69 0.13
VB1 0.15 0.00 0.02 0.16 0.00 0.03
VB2 0.16 0.32 0.02 0.18 0.53 0.02
ADVB 0.22 0.26 0.08 0.22 0.10 0.07
Table 2 provides a summary of the analysis results. As the results show, our proposed method produces an estimated value function similar to the estimated optimal policy, while being considerably fairer in terms of both action fairness and outcome fairness.
7 Discussion
This work studies double fairness in policy learning by simultaneously addressing three objectives. We provide a systematic analysis to identify conditions under which a outcome-fair or double-fair policy can exist. We also propose a novel estimation framework that incorporates fairness constraints directly into the multi-objective policy learning problem. Our approach is flexible and accommodates a range of fairness notions. Extensive experiments on both synthetic and real-world datasets validate our method, showcasing its effectiveness in achieving double fairness without compromising overall welfare.
Supplemental Materials for “Doubly Trustworthy Policy Learning”
8 Additional Results
8.1 Limitation of the Linear Scalarization Approach
In this subsection, we show that the limitation of the linear scalarization approach via the following simple example.
(a) Linear scalarization.
(b) Weighted Tchebycheff scalarization.
Figure 7: Linear scalarization approach cannot identify Pareto polices in the non-convex regions of the Pareto front, as these solutions lack supporting hyperplanes (Miettinen, 1999). In contrast, weighted Tchebycheff scalarization can find all Pareto policies, regardless of the shape of the objective space. Figure adapted from Lin et al. (2024).
Consider the following data generation procedure: $S\sim\mbox{Ber}(0.5)$, $X\sim\mbox{Ber}(0.6)$, R=S+X+0.4A0.5ASR=S+X+0.4A-0.5AS. The policy class $\Pi$ consists the following three policies: $\pi_{1}=\mathds{1}(X>0)$, $\pi_{2}=\mathds{1}(S+X>0)$, $\pi_{3}=\mathds{1}(S>0)$. That is, $\Pi=\{\pi_{1},\pi_{2},\pi_{3}\}$.
Figure 8: The Pareto front.
Under this data generation process, it can be easily shown that
$\Delta_{1}(\pi_{1})=0,\Delta_{2}(\pi_{1})=0.7,$
$\Delta_{1}(\pi_{2})=0.4,\Delta_{2}(\pi_{2})=0.66,$
$\Delta_{1}(\pi_{3})=1,\Delta_{2}(\pi_{3})=0.5,$
for the equal opportunity fairness. See Figure 8 below for the Pareto front. Clearly, all the three polices lies in the Pareto set, i.e., they cannot Pareto dominate each other.
Using the linear scalarization approach, it can be shown that
$\left\{\pi\in\Pi:\operatorname*{arg\,min}_{\pi}\alpha\Delta_{1}(\pi)+(1-\alpha)\Delta_{2}(\pi),\;\forall\alpha\in(0,1)\right\}=\{\pi_{1},\pi_{3}\}.$
That is, there does not exist a $\alpha\in(0,1)$ such that $\alpha\Delta_{1}(\pi_{2})+(1-\alpha)\Delta_{2}(\pi_{2})<\alpha\Delta_{1}(\pi_{1})+(1-\alpha)\Delta_{2}(\pi_{1})$, and $\alpha\Delta_{1}(\pi_{2})+(1-\alpha)\Delta_{2}(\pi_{2})<\alpha\Delta_{1}(\pi_{3})+(1-\alpha)\Delta_{2}(\pi_{3})$. As a result, the linear scalarization can only identify $\pi_{1}$ and $\pi_{3}$, while leaving the policy that can yield the highest value-$\pi_{2}$ out of the Pareto set.
In comparison, the lexicographic weighted Tchebyshev scalarization can successfully identify all the three policies. To show this, only basic algebraic calculations are needed; hence, the procedure is omitted for brevity.
8.2 Existence of Counterfactual outcome fairness Policy
In this subsection, we briefly discuss how the results in Propositions 1 and 2 can be extended to counterfactual fairness. For brevity, we omit the proof, as it can be derived in a similar manner to the proofs of Propositions 1 and 2.
Assumption 8.1.
Given the subject with information $x$, for any $a\neq a^{\prime}$, there exists a treatment option $a^{\prime}$ dominates the other treatment choice $a$, regardless of the sensitive group, i.e., $\min_{s}f(s,x,a^{\prime})\geq\max_{s}f(s,x^{\prime},a)$.
Assumption 8.2.
(i) $f(1,x(1),1)-f(0,x(0),1)][f(1,x(1),0)-f(0,x(0),0)]\leq 0$, for counterfactual covariates $x(1)$ and $x(0)$, almost surely.
(ii) $f(1,x(1),1)-f(1,x(1),0)][f(0,x(0),1)-f(0,x(0),0)]\leq 0$, for counterfactual covariates $x(1)$ and $x(0)$, almost surely.
(iii) Given subject information $x$, there exists a pair $(s,a)$ such that $f(s,x,a)\geq f(s,x,a^{\prime})$, and $f(s,x,a)\geq f(s^{\prime},x^{\prime},a)$, for all $x^{\prime}$, almost surely.
The following Propositions 4 and 5 establish the sufficient and necessary conditions for the existence of the outcome fair policy and the double fairness policy, respectively.
Proposition 4.
There exists a counterfactual outcome fair policy for the environment if and only if either Assumption 8.1 or Assumption 8.2.
Proposition 5.
A double counterfactual fairness policy is is either unique or does not exist. Furthermore, a unique double fairness policy exists if and only if Assumption 8.2 (i) holds.
8.3 Extension to Discrete Action Space
In this subsection, we briefly discuss how the results in Propositions 1 and 2 can be extended to discrete action spaces. For brevity, we only discuss the extension for the equal opportunity fairness, while the counterfactual fairness can be generalized analogously.
Throughout this subsection, we assume the action space is $L$-dimensional, and $\mathcal{A}=\{1,2,\dots,L\}.$ Let $\bm{f}(s,x)=\left[f(s,x,1),f(s,x,2),\dots,f(s,x,L)\right]^{\top},$ and $\bm{\pi}(s,x)=\left[\pi(1|s,x),\dots,\pi(L|s,x)\right]^{\top},$ for any $s$ and $x$. It then follows that for equal opportunity, we have
$\bm{f}(s,x)^{\top}\bm{\pi}(s,x)=\bm{f}(s^{\prime},x)^{\top}\bm{\pi}(s^{\prime},x),$
for any $s$ and $x$. Thus, a outcome fair policy exists, if and only if the above linear equation has solutions, subject to $\bm{\pi}^{\top}\bm{1}_{L}=1$, where $\bm{1}_{L}$ is all-ones vector with dimension $L$; and $0\leq\pi(l|s,x)\leq 1$, for any $l=1,2,\dots,L$.
Similarly, a double fairness policy exist if and only if the above linear equation has solutions, subject to $\bm{\pi}^{\top}\bm{1}_{L}=1$, and $0\leq\pi(l|s,x)=\pi(l|s^{\prime},x)\leq 1$, for any $l=1,2,\dots,L$.
8.4 Further Details on Premium in the Insurance Data
In the offline dataset, we classify customers as low-risk or high-risk based on the BM scale, following common practice in the insurance literature, where drivers are often categorized as good or bad based on their BM score (Frangos and Vrontos, 2001). The BM scale ranges from 0 to 22, with higher values indicating greater risk. We define customers with BM greater than five as high-risk and the rest as low-risk. Since the dataset does not contain premium information, we generate the pure premium amount using the frequency-severity approach, as outlined by Baumann and Loi (2023). This approach sets pure premiums based on the predicted risk. To better reflect real-world pricing, we further adjust the final premium based on the coverage plan and gender. Specifically, the premium is generated according to
$\min\left(500\times\texttt{coverage}+100\times\frac{\#\texttt{claims}}{\texttt{exposure}}+20\times\texttt{gender},\ 5000\right),$
where coverage is a categorical variable representing one of three coverage plans, and gender is a binary variable (e.g., 0 for female, 1 for male). The term #claims refers to the number of insurance claims made by the policyholder, and exposure denotes the duration that the policy was active. The final reward for a subject is defined as the difference between the adjusted premium and the total claim amount.
9 Proofs and Auxiliary Lemmas
Throughout, we use $c$ and $C$ to denote a generic constant that can vary from line to line.
9.1 Proof of Proposition 1
Proof.
It suffices to solve the following linear equation with respect to $\pi$:
$f(0,x,1)\pi(1|0,x)+f(0,x,0)\pi(0|0,x)=f(1,x,1)\pi(1|1,x)+f(1,x,0)\pi(0|1,x).$
However, the above linear systems have two unknowns and only one equation, thus it could have infinity many solutions. Denote $\pi(0,x)$ as $\pi(1|S=0,x)$, this yields that
$\pi(1,x)=\left[(f(0,x,1)-f(0,x,0))\pi(0,x)+f(0,x,0)-f(1,x,0)\right][f(1,x,1)-f(1,x,0)]^{-1}.$ (5)
It is equivalent to show that there exists a $\pi(0,x)\in[0,1]$ such that its corresponding $\pi(1,x)\in[0,1]$.
By Equation (5), $\pi(1,x)$ is linear in $\pi(0,x)$. Thus, we have either
$\pi(1,x)\in\left[\frac{f(0,x,0)-f(1,x,0)}{f(1,x,1)-f(1,x,0)},\frac{f(0,x,1)-f(1,x,0)}{f(1,x,1)-f(1,x,0)}\right],$ (6)
$\mbox{or }\pi(1,x)\in\left[\frac{f(0,x,1)-f(1,x,0)}{f(1,x,1)-f(1,x,0)},\frac{f(0,x,0)-f(1,x,0)}{f(1,x,1)-f(1,x,0)}\right].$ (7)
Thus it remains to show that either the above two intervals intersects with the interval $[0,1]$ is non-empty. By some algebra, interval (6) has intersection with $[0,1]$ if and only if the following holds for any $a\neq a^{\prime}$,
$\begin{cases}f(1,x,a)>f(1,x,a^{\prime}),\\ f(1,x,a)>f(0,x,a^{\prime}),\\ f(0,x,a)>f(1,x,a^{\prime}),\\ f(0,x,a)>f(0,x,a^{\prime}).\end{cases}$
This corresponds to Assumption 3.1. Following the same argument, interval (7) has intersection with $[0,1]$ if and only if for $s\neq s^{\prime}$ and $a\neq a^{\prime}$,
$\begin{cases}f(s,x,a)>f(s,x,a^{\prime}),\\ f(s,x,a)>f(s^{\prime},x,a),\\ f(s^{\prime},x,a^{\prime})>f(s^{\prime},x,a),\\ f(s^{\prime},x,a^{\prime})>f(s,x,a^{\prime}),\end{cases}$
which corresponds to Assumption 3.2. This thus completes the proof. ∎
9.2 Proof of Proposition 2
Proof.
Following the proof of Proposition 1, by letting $\pi(1,x)=\pi(0,x)=\pi(x)$, we obtain that
$\pi(x)=[f(0,x,0)-f(1,x,0)]\left[f(0,x,0)-f(1,x,0)+f(1,x,1)-f(0,x,1)\right]^{-1}.$ (8)
Thus, if $\pi(x)\in[0,1]$, then the solution is unique; and if $\pi(x)\notin[0,1]$, there is no double fairness policy exists.
By (8), $\pi(x)\in(0,1)$ if and only if either one of the following two scenarios holds.
  1. 1.
    f(1,x,1)f(0,x,1)<0f(1,x,1)-f(0,x,1)<0 and f(0,x,0)f(1,x,0)<0f(0,x,0)-f(1,x,0)<0;
  2. 2.
    f(1,x,1)f(0,x,1)>0f(1,x,1)-f(0,x,1)>0 and f(0,x,0)f(1,x,0)>0f(0,x,0)-f(1,x,0)>0.
The above two equations correspond to Assumption 3.2 (i). This thus completes the proof.
9.3 Proof of Proposition 3
Proof.
To show all of the elements in $\Pi_{p}$ is Pareto optimal solution, see Theorem 5.4.1 in Miettinen (1999); while the other direction can be proved using Theorem 5.4.2 in Miettinen (1999). ∎
9.4 Proof of Theorem 1
Lemma 1.
Suppose that the assumptions in Theorem 1 hold, then for any $\alpha\in(0,1)$, there exists some $k\in\left\{1,2,\cdots,K\right\}$ such that the following holds almost surely:
$\left|M_{\alpha}^{*}-\widehat{M}_{\alpha_{k}}^{*}\right|\lesssim\sup_{\pi\in\Pi}\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|+\sup_{\pi\in\Pi}\left|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\right|+K^{-1},$
$\mbox{ and }\left|\Delta_{\alpha}^{*}-\widehat{\Delta}_{\alpha_{k}}^{*}\right|\lesssim\sup_{\pi\in\Pi}\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|+\sup_{\pi\in\Pi}\left|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\right|+K^{-1}.$
Proof.
For any $\alpha\in(0,1)$, there exists $\alpha_{k}$ such that $\alpha_{k}\in\Omega$ and $|\alpha_{k}-\alpha|\leq K^{-1}$. We now first prove the bound for $\big|M_{\alpha}^{*}-\widehat{M}_{\alpha_{k}}^{*}\big|$. By definition, we have
$M_{\alpha}^{*}=\inf_{\pi\in\Pi}\max\left\{\alpha\Delta_{1}(\pi),(1-\alpha)\Delta_{2}(\pi)\right\},$
$\widehat{M}_{\alpha_{k}}^{*}=\inf_{\pi\in\Pi}\max\big\{\alpha_{k}\widehat{\Delta}_{1}(\pi),(1-\alpha_{k})\widehat{\Delta}_{2}(\pi)\big\}.$
It then follows that
$\left|M_{\alpha}^{*}-\widehat{M}_{\alpha_{k}}^{*}\right|$
$=$ $\left|\inf_{\pi\in\Pi}\left[\max\left(\alpha\Delta_{1}(\pi),(1-\alpha)\Delta_{2}(\pi)\right)\right]-\inf_{\pi\in\Pi}\left[\max\left(\alpha_{k}\widehat{\Delta}_{1}(\pi),(1-\alpha_{k})\widehat{\Delta}_{2}(\pi)\right)\right]\right|$
$\leq$ $\;\sup_{\pi\in\Pi}\left|\max\left(\alpha\Delta_{1}(\pi),(1-\alpha)\Delta_{2}(\pi)\right)-\max\left(\alpha_{k}\widehat{\Delta}_{1}(\pi),(1-\alpha_{k})\widehat{\Delta}_{2}(\pi)\right)\right|$ (9)
$\leq$ $\;\sup_{\pi\in\Pi}\max\Big[\left|\alpha\Delta_{1}(\pi)-\alpha_{k}\widehat{\Delta}_{1}(\pi)\right|,\left|(1-\alpha)\Delta_{2}(\pi)-(1-\alpha_{k})\widehat{\Delta}_{2}(\pi)\right|\Bigg]$
$\leq$ $\;\sup_{\pi\in\Pi}\left|\alpha\Delta_{1}(\pi)-\alpha_{k}\widehat{\Delta}_{1}(\pi)\right|+\sup_{\pi\in\Pi}\left|(1-\alpha)\Delta_{2}(\pi)-(1-\alpha_{k})\widehat{\Delta}_{2}(\pi)\right|,$
where the second to last inequality follows from triangle inequality of the $\ell_{\infty}$ norm.
We first deal with term $\sup_{\pi\in\Pi}\left|\alpha\Delta_{1}(\pi)-\alpha_{k}\widehat{\Delta}_{1}(\pi)\right|$, the second term can be bounded similarly. It follows that
$\sup_{\pi\in\Pi}\left|\alpha\Delta_{1}(\pi)-\alpha_{k}\widehat{\Delta}_{1}(\pi)\right|=\sup_{\pi\in\Pi}\left|\alpha\Delta_{1}(\pi)-\alpha_{k}\Delta_{1}(\pi)+\alpha_{k}\Delta_{1}(\pi)-\alpha_{k}\widehat{\Delta}_{1}(\pi)\right|$
$\leq$ $\;\underbrace{\sup_{\pi\in\Pi}\left|\alpha\Delta_{1}(\pi)-\alpha_{k}\Delta_{1}(\pi)\right|}_{J_{1}}+\underbrace{\sup_{\pi\in\Pi}\left|\alpha_{k}\Delta_{1}(\pi)-\alpha_{k}\widehat{\Delta}_{1}(\pi)\right|}_{J_{2}}.$ (10)
Since $\alpha_{k}\in\Omega$, the first term $J_{1}$ is upper bounded by $K^{-1}$, by the fact that $\pi(a|s,x)\in[0,1]$ for any $a$, $s$, and $x$, and thus $\sup_{\pi}|\Delta_{1}(\pi)|\leq 1$. The second term $J_{2}$ is simply bounded by $\sup_{\pi\in\Pi}\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|$. Thus we have proved that
$\sup_{\pi\in\Pi}\left|\alpha\Delta_{1}(\pi)-\alpha_{k}\widehat{\Delta}_{1}(\pi)\right|\leq K^{-1}+\sup_{\pi\in\Pi}\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|.$
Using the similar argument, we can show that
$\sup_{\pi\in\Pi}\left|(1-\alpha)\Delta_{2}(\pi)-(1-\alpha_{k})\widehat{\Delta}_{2}(\pi)\right|\lesssim K^{-1}+\sup_{\pi\in\Pi}\left|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\right|.$
Finally, the term $\left|\Delta_{\alpha}^{*}-\widehat{\Delta}_{\alpha_{k}}^{*}\right|$ can be bounded analogously and we omit the details. This completes the proof. ∎
We now prove Theorem 1.
Proof.
We first show if the slack parameter $\kappa$ is chosen such that
$\kappa\geq C\max\left[\sup_{\pi\in\Pi}\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|+\sup_{\pi\in\Pi}\left|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\right|,\;K^{-1}\right],$
for a sufficiently large constant $C$, then the estimated Pareto set covers the underlying true Pareto set, almost surely.
Recall that
$\Pi_{p}=\bigcup_{\alpha\in(0,1)}\left\{\pi\in\Lambda_{\alpha}^{*}:\Delta(\pi)\leq\Delta_{\alpha}^{*}\right\},$
 $\mbox{ and }\widehat{\Pi}_{p}=\bigcup_{k=1}^{K}\left\{\pi\in\widehat{\Lambda}_{\alpha_{k}}^{*}:\widehat{\Delta}(\pi)\leq\widehat{\Delta}_{\alpha_{k}}^{*}+\kappa\right\}.$
It is therefore equivalent to show that, for any $\alpha\in(0,1)$ and $\pi$:
$\Delta(\pi)\leq\Delta_{\alpha}^{*}\mbox{ and }M_{\alpha}(\pi)=M_{\alpha}^{*},$
$\implies\exists\,\alpha_{k},\widehat{\Delta}(\pi)\leq\widehat{\Delta}_{\alpha_{k}}^{*}+\kappa\mbox{ and }\widehat{M}_{\alpha_{k}}(\pi)\leq\widehat{M}_{\alpha_{k}}^{*}+\kappa.$
It thus suffices to show that, for any $\alpha\in(0,1)$ and $\pi$:
$M_{\alpha}(\pi)=M_{\alpha}^{*}\implies\widehat{M}_{\alpha_{k}}(\pi)\leq\widehat{M}_{\alpha_{k}}^{*}+\kappa,$ (11)
 $\mbox{ and }\Delta(\pi)\leq\Delta_{\alpha}^{*}\implies\exists\,\alpha_{k},\widehat{\Delta}(\pi)\leq\widehat{\Delta}_{\alpha_{k}}^{*}+\kappa.$
We first show that $M_{\alpha}(\pi)=M_{\alpha}^{*}\implies\widehat{M}_{\alpha_{k}}(\pi)\leq\widehat{M}_{\alpha_{k}}^{*}+\kappa.$ Recall that
$\widehat{M}_{\alpha_{k}}(\pi)=\max\left(\alpha_{k}\widehat{\Delta}_{1}(\pi),(1-\alpha_{k})\widehat{\Delta}_{2}(\pi)\right)\mbox{ and }\widehat{M}_{\alpha_{k}}^{*}=\inf_{\pi\in\Pi}\widehat{M}_{\alpha_{k}}(\pi).$
It follows that
$\widehat{M}_{\alpha_{k}}(\pi)=\max\left(\alpha_{k}\widehat{\Delta}_{1}(\pi),(1-\alpha_{k})\widehat{\Delta}_{2}(\pi)\right)$
$=$ $\max\left(\alpha_{k}\widehat{\Delta}_{1}(\pi),(1-\alpha_{k})\widehat{\Delta}_{2}(\pi)\right)-\max\left(\alpha\Delta_{1}(\pi),(1-\alpha)\Delta_{2}(\pi)\right)$
$+\max\left(\alpha\Delta_{1}(\pi),(1-\alpha)\Delta_{2}(\pi)\right)$
$\lesssim$ $\sup_{\pi\in\Pi}\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|+\sup_{\pi\in\Pi}\left|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\right|+K^{-1}+\max\left(\alpha\Delta_{1}(\pi),(1-\alpha)\Delta_{2}(\pi)\right)$
[…]
$\lesssim$ $\;\sup_{\pi\in\Pi}\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|+\sup_{\pi\in\Pi}\left|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\right|+K^{-1}+M_{\alpha}^{*}-\widehat{M}_{\alpha_{k}}^{*}+\widehat{M}_{\alpha_{k}}^{*}$
[…]
Since $\kappa$ satisfies
$\kappa\geq C\max\left[\sup_{\pi\in\Pi}\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|+\sup_{\pi\in\Pi}\left|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\right|,\;K^{-1}\right],$
for a sufficiently large constant $C$, we have proved Equation (11). Another argument that $\Delta(\pi)\leq\Delta_{\alpha}^{*}\implies\exists\,\alpha_{k},\;\widehat{\Delta}(\pi)\leq\widehat{\Delta}_{\alpha_{k}}^{*}+\kappa$ can be proved using the similarly reasoning, we omit the details to save the space.
[…]
where $\widehat{V}^{*}$ is the value estimator of $V^{*}$ using the reward estimator. The term $I_{1}\leq 0$ since $\widehat{V}(\widehat{\pi})=\sup_{\pi\in\widehat{\Pi}_{p}}\widehat{V}(\pi)$; and by Theorem 2, the estimated Pareto set contains the optimal policy in the Pareto set, i.e., $\pi^{*}\in\widehat{\Pi}_{p}$.
It then follows that
[…]
$\leq$ $\;\sup_{\pi\in\Pi_{p}}|V({\pi})-\widehat{V}({\pi})|+\sup_{\pi\in\widehat{\Pi}_{p}}|V({\pi})-\widehat{V}({\pi})|.$
The above term can be upper bounded by $2\sup_{\pi\in\widehat{\Pi}_{p}}|V({\pi})-\widehat{V}({\pi})|$, and hence further upper bounded by
$2\sup_{\pi\in\Pi}|V({\pi})-\widehat{V}({\pi})|.$
Another upper bound can be established in the following way, and we use Hausdorff distance to “match” each $\pi\in\widehat{\Pi}_{p}$ with a nearby $\pi^{\prime}\in\Pi_{p}$. Recall that
$\eta(\pi)\equiv\widehat{V}(\pi)-V(\pi)$
denotes the policy evaluation error. We have
$\sup_{\pi\in\widehat{\Pi}_{p}}|V({\pi})-\widehat{V}({\pi})|=\sup_{\pi\in\widehat{\Pi}_{p}}|\eta(\pi)|$
$=$ $\sup_{\pi\in\widehat{\Pi}_{p}}\sup_{\begin{subarray}{c}\pi^{\prime}\in\Pi_{p}:\\ d_{\mathrm{TV}}(\pi,\pi^{\prime})\leq H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})\end{subarray}}|\eta(\pi)-\eta(\pi^{\prime})+\eta(\pi^{\prime})|$
$\leq$ $\;\sup_{d_{\mathrm{TV}}(\pi,\pi^{\prime})\leq H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})}\left|\eta({\pi})-\eta({\pi^{\prime}})\right|+\sup_{\pi\in\Pi_{p}}|\eta({\pi})|$
$\leq$ $\;\sup_{d_{\mathrm{TV}}(\pi,\pi^{\prime})\leq H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})}|V({\pi})-V({\pi^{\prime}})|+\sup_{\pi\in\Pi_{p}}|\eta({\pi})|$
$+\sup_{d_{\mathrm{TV}}(\pi,\pi^{\prime})\leq H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})}|\widehat{V}({\pi})-\widehat{V}({\pi^{\prime}})|$
$\leq$ $\;\sup_{d_{\mathrm{TV}}(\pi,\pi^{\prime})\leq H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})}\left|\mathbb{E}\;\sum_{a}r(S,X,a)(\pi(a|S,X)-\pi^{\prime}(a|S,X))\right|+\sup_{\pi\in\Pi_{p}}|\eta({\pi})|$
$+\sup_{d_{\mathrm{TV}}(\pi,\pi^{\prime})\leq H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})}\left|\frac{1}{n}\sum_{i=1}^{n}\sum_{a}\widehat{r}(S_{i},X_{i},a)(\pi(a|S_{i},X_{i})-\pi^{\prime}(a|S_{i},X_{i}))\right|.$
We first bound the first term, and the last term can be bounded following the similar logic. The first term
$\sup_{d_{\mathrm{TV}}(\pi,\pi^{\prime})\leq H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})}\left|\mathbb{E}\;\sum_{a}r(S,X,a)(\pi(a|S,X)-\pi^{\prime}(a|S,X))\right|$
$\leq\;$ $\sup_{d_{\mathrm{TV}}(\pi,\pi^{\prime})\leq H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})}\sup_{s,x}\sum_{a}\left|r(s,x,a)(\pi(a|s,x)-\pi^{\prime}(a|s,x))\right|$
$\leq\;$ $\sup_{d_{\mathrm{TV}}(\pi,\pi^{\prime})\leq H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})}\underbrace{\sup_{s,x}\sum_{a}|\pi(a|s,x)-\pi^{\prime}(a|s,x)|}_{2d_{\mathrm{TV}}(\pi,\pi^{\prime})}\;R_{\max}\mbox{ (Hölder's inequality )}$
$\leq\;$ $2R_{\max}H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})$
Similarly, we have
$\left|\frac{1}{n}\sum_{i=1}^{n}\sum_{a}\widehat{r}(S_{i},X_{i},a)(\pi(a|S_{i},X_{i})-\pi^{\prime}(a|S_{i},X_{i}))\right|\leq 2R_{\max}H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p}).$
To summarize, we have provided the upper bound for the regret, which is
$\min\left[\sup_{\pi\in\Pi_{p}}\big|\eta(\pi)\big|+H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p}),\;\sup_{\pi\in\Pi}\big|\eta(\pi)\big|\right],$
up to a constant factor.
We next prove the other direction.
[…]
$\geq$ $\;\sup_{\pi\in\Pi_{p}}V({\pi})-\sup_{\pi\in\widehat{\Pi}_{p}}V({\pi})$
$\geq$ $\;-2R_{\max}H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p}),$
where the last inequality follows from
$\sup_{\pi\in\widehat{\Pi}_{p}}V({\pi})\leq\sup_{\pi\in\Pi_{p}}V({\pi})+2R_{\max}H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p}),$
analogous to the argument we used earlier for the upper bound.
This completes the proof. ∎
9.5 Proof of Theorem 2
Proof.
In the proof of Theorem 1, we have demonstrated that the estimated Pareto set covers the underlying true Pareto set with an appropriate choice of $\kappa$. We next provide the upper bound of the Hausdorff distance $H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})$. Since $\Pi_{p}\subseteq\widehat{\Pi}_{p}$, we have for every $\pi\in\Pi_{p}$, $d_{\mathrm{TV}}(\pi,\widehat{\Pi}_{p})=0$ (take $\pi$ itself in $\widehat{\Pi}_{p}$), and hence $\sup_{\pi\in\Pi_{p}}d_{\mathrm{TV}}(\pi,\widehat{\Pi}_{p})=0$. Therefore,
$H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})=\max\big[\sup_{\pi\in\Pi_{p}}d_{\mathrm{TV}}(\pi,\widehat{\Pi}_{p}),\sup_{\pi\in\widehat{\Pi}_{p}}d_{\mathrm{TV}}(\pi,\Pi_{p})\big]=\sup_{\pi\in\widehat{\Pi}_{p}}d_{\mathrm{TV}}(\pi,\Pi_{p}).$
Because $\widehat{\Pi}_{p}\subseteq\Pi$, by Assumption 4.2, to show that $\sup_{\pi\in\widehat{\Pi}_{p}}d_{\mathrm{TV}}(\pi,\Pi_{p})\leq\delta$, where $\delta$ represents the target bound, it suffices to prove that $\forall\;\widetilde{\pi}\in\widehat{\Pi}_{p}$, $\exists\;\alpha(\widetilde{\pi})\in(0,1)$, such that
$\left|\Delta(\widetilde{\pi})-\Delta_{\alpha(\widetilde{\pi})}^{*}\right|\leq c\delta,\mbox{ and }\left|M_{\alpha(\widetilde{\pi})}(\widetilde{\pi})-M_{\alpha}^{*}\right|\leq c\delta,$
where $c>0$ is the universal constant in Assumption 4.2.
We first show that $\forall\;\widetilde{\pi}\in\widehat{\Pi}_{p}$, $\exists\;\alpha(\widetilde{\pi})\in(0,1)$, such that $\left|M_{\alpha(\widetilde{\pi})}(\widetilde{\pi})-M_{\alpha(\widetilde{\pi})}^{*}\right|\leq c\delta$. It follows that
[…]
$\leq\;$ $\underbrace{\left|M_{\alpha(\widetilde{\pi})}(\widetilde{\pi})-\widehat{M}_{\alpha(\widetilde{\pi})}(\widetilde{\pi})\right|}_{Q_{1}}+\underbrace{\left|\widehat{M}_{\alpha(\widetilde{\pi})}(\widetilde{\pi})-M_{\alpha(\widetilde{\pi})}^{*}\right|}_{Q_{2}}.$
We first bound $Q_{2}$. Recall that by Equation (2), we have
$\widehat{\Pi}_{p}=\bigcup_{k=1}^{K}\left\{\pi\in\widehat{\Lambda}_{\alpha_{k}}^{*}:\widehat{\Delta}(\pi)\leq\widehat{\Delta}_{\alpha_{k}}^{*}+\kappa\right\}\text{ and }\widehat{\Lambda}_{\alpha}^{*}=\left\{\pi\in\Pi:\widehat{M}_{\alpha}(\pi)\leq\widehat{M}_{\alpha}^{*}+\kappa\right\}.$
[…]
$\leq$ $\kappa+\sup_{\pi\in\Pi}\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|+\sup_{\pi\in\Pi}\left|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\right|,$
[…]
It remains to bound term $Q_{1}$. It follows that
[…]
$\leq$ $\;\sup_{\alpha\in(0,1)}\sup_{\pi\in\Pi}\Big|\max\left(\alpha\Delta_{1}(\pi),(1-\alpha)\Delta_{2}(\pi)\right)-\max\left(\alpha\widehat{\Delta}_{1}(\pi),(1-\alpha)\widehat{\Delta}_{2}(\pi)\right)\Big|$
$\leq$ $\;\sup_{\alpha\in(0,1)}\sup_{\pi\in\Pi}\Big|\max\Big[\alpha\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|,(1-\alpha)\left|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\right|\Big]\Big|$
$\leq$ $\;\sup_{\pi\in\Pi}\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|+\sup_{\pi\in\Pi}\left|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\right|.$
Thus, we have $\forall\;\widetilde{\pi}\in\widehat{\Pi}_{p}$, $\exists\;\alpha(\widetilde{\pi})\in(0,1)$, such that
[…]
$\lesssim$ $\;\sup_{\pi\in\Pi}\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|+\sup_{\pi\in\Pi}\left|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\right|+\kappa$
Similarly, we can show that $\forall\;\widetilde{\pi}\in\widehat{\Pi}_{p}$, $\exists\;\alpha(\widetilde{\pi})\in(0,1)$, such that
$\left|\Delta(\widetilde{\pi})-\Delta_{\alpha}^{*}\right|\lesssim\sup_{\pi\in\Pi}\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|+\sup_{\pi\in\Pi}\left|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\right|+\kappa.$
Therefore, by Assumption 4.2, we have
$H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})=\sup_{\pi\in\widehat{\Pi}_{p}}d_{\mathrm{TV}}(\pi,\Pi_{p})$
$\lesssim\;$ $\sup_{\pi\in\Pi}\left|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\right|+\sup_{\pi\in\Pi}\left|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\right|+\kappa.$
Finally, recall that
$\kappa\geq C\max\Big[\sup_{\pi\in\Pi}\big|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\big|+\sup_{\pi\in\Pi}\big|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\big|,\;K^{-1}\Big],$
for some constant $C$. Thus, we have
$H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})\lesssim\kappa.$
This completes the proof. ∎
9.6 Proof of Theorem 3: Equal Opportunity
Proof.
By Theorems 1 and 2, it remains to
  1. 1.
    Establish an upper bound for $\sup_{\pi\in\Pi}\big|\widehat{V}(\pi)-V(\pi)\big|$;
  2. 2.
    Establish an upper bound for
    $\sup_{\pi\in\Pi}\big|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\big|+\sup_{\pi\in\Pi}\big|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\big|,$
    under the equal opportunity fairness notion.
We now present the following lemmas.
Lemma 2 (Uniform Convergence of the Value Estimator).
$\sup_{\pi\in\Pi}|\widehat{V}(\pi)-V(\pi)|=O_{p}\left(\sqrt{\mbox{VC}(\Pi)/n}+n^{-\gamma_{1}/2}+\sqrt{\log n/n}\right).$
Proof.
$\sup_{\pi\in\Pi}|\widehat{V}(\pi)-V(\pi)|$
$=$ $\;\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum_{i=1}^{n}\sum_{a}\widehat{r}(S_{i},X_{i},a)\pi(a|S_{i},X_{i})-\mathbb{E}\sum_{a}r(S,X,a)\pi(a|S,X)\right|$
$\leq$ $\;\sup_{\pi\in\Pi}\left|\sum_{a}\frac{1}{n}\sum_{i=1}^{n}(\widehat{r}(S_{i},X_{i},a)-r(S_{i},X_{i},a))\pi(a|S_{i},X_{i})\right|$
$+$ $\;\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum_{i=1}^{n}r(S_{i},X_{i},a)\pi(a|S_{i},X_{i})-\mathbb{E}\sum_{a}r(S,X,a)\pi(a|S,X)\right|.$
Applying the Cauchy-Schwarz, the first term above is upper bounded by
$\sqrt{\max_{a}\frac{1}{n}\sum_{i=1}^{n}(\widehat{r}(S_{i},X_{i},a)-r(S_{i},X_{i},a))^{2}}=O_{p}\left(n^{-\frac{\gamma_{1}}{2}}\right),$
by Assumption 4.4. Using standard empirical process techniques (see, e.g., Van Der Vaart and Wellner (1996)), the second term is of order
$O_{p}\left(\sqrt{\frac{\mathrm{VC}(\Pi)}{n}}+\sqrt{\frac{\log n}{n}}\right).$
This completes the proof. ∎
Lemma 3 (Equal Opportunity Action Fairness Evaluation).
For the equal opportunity fair policy, we have
$\sup_{\pi\in\Pi}|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)|=O_{p}\left(\sqrt{\mbox{VC}(\Pi)/n}+\sqrt{\log n/n}\right)$
Proof.
It suffices to show that
$\mathbb{E}\sup_{\pi\in\Pi}|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)|=O\left(\sqrt{\mbox{VC}(\Pi)/n}\right).$
Let $\Delta_{1i}\equiv\left[\pi(1,X_{i})-\pi(0,X_{i})\right]^{2}$, and $\{\epsilon_{i}\}_{i}$ be i.i.d. Rademacher random variables. Because $\Delta_{1}(\pi)$ is an expectation of i.i.d. copies of $\Delta_{1i}$ and $\widehat{\Delta}_{1}(\pi)$ is the corresponding empirical mean, we obtain by applying the symmetrization technique,
$\mathbb{E}\sup_{\pi\in\Pi}|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)|\leq 2\mathbb{E}\sup_{\pi\in\Pi}\left\lvert\cfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\Delta_{1i}(\pi)\right\rvert$
$=2\mathbb{E}\sup_{\pi\in\Pi}\cfrac{1}{n}\left\lvert\sum_{i=1}^{n}\epsilon_{i}(\pi(1,X_{i})-\pi(0,X_{i}))^{2}\right\rvert$
$\leq 2\mathbb{E}\sup_{\pi\in\Pi}\cfrac{1}{n}\left\lvert\sum_{i=1}^{n}\epsilon_{i}\pi^{2}(1,X_{i})\right\rvert+2\mathbb{E}\sup_{\pi\in\Pi}\cfrac{1}{n}\left\lvert\sum_{i=1}^{n}\epsilon_{i}\pi^{2}(0,X_{i})\right\rvert$
$+4\mathbb{E}\sup_{\pi\in\Pi}\cfrac{1}{n}\left\lvert\sum_{i=1}^{n}\epsilon_{i}\pi(0,X_{i})\pi(1,X_{i})\right\rvert$.
Since $\Pi$ is a VC-subgraph class and $\pi(\cdot,x)\in[0,1]$, the transformed classes
$\{\pi^{2}(1,\cdot):\pi\in\Pi\},\qquad\{\pi^{2}(0,\cdot):\pi\in\Pi\},\qquad\{\pi(1,\cdot)\pi(0,\cdot):\pi\in\Pi\}$
are still VC-type, because squaring is a monotone Lipschitz map on $[0,1]$ and products of bounded VC-type classes remain VC-type (see Lemma 2.6.20 in Van Der Vaart and Wellner (1996)). Thus, a direct application of Dudley’s entropy integral bound yields that
$\mathbb{E}\sup_{\pi\in\Pi}\left\lvert\cfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\Delta_{1i}(\pi)\right\rvert=O(\sqrt{\mbox{VC}(\Pi)/n}).$
Define $Z(X_{1},\ldots,X_{n}):=\sup_{\pi\in\Pi}\bigl|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\bigr|$ Fix an index $i$ and let $\mathbf{X}=(X_{1},\ldots,X_{n})$ and $\mathbf{X}^{(i)}=(X_{1},\ldots,X_{i}^{\prime},\ldots,X_{n})$ differ only at coordinate $i$. We have $|Z(\mathbf{X})-Z(\mathbf{X}^{(i)})|\leq n^{-1}$. By bounded difference inequality, for any $t>0$,
$\Pr\bigl(Z-\mathbb{E}Z\geq t\bigr)\leq\exp(-2nt^{2}),$
Taking $t=\sqrt{\log n/(2n)}$ gives
$Z\leq\mathbb{E}Z+\sqrt{\frac{\log n}{2n}}\quad\text{with probability at least }1-\frac{1}{n}.$
Combining with $\mathbb{E}Z=O(\sqrt{\mbox{VC}(\Pi)/n})$ yields
$\sup_{\pi\in\Pi}|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)|=O_{p}\left(\sqrt{\mbox{VC}(\Pi)/n}+\sqrt{\log n/n}\right).$
This completes the proof. ∎
Lemma 4 (Equal Opportunity Outcome fairness Evaluation).
For the equal opportunity fair policy, the following holds:
$\sup_{\pi\in\Pi}|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)|=O_{p}\left(\sqrt{\mbox{VC}(\Pi)/n}+n^{-\gamma_{2}/2}+\sqrt{\log n/n}\right).$
Proof.
By the definition, we have
$\sup_{\pi\in\Pi}|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)|$
$=$ $\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}^{\pi}(1,X_{i})-\widehat{f}^{\pi}(0,X_{i}))^{2}-\mathbb{E}\left[(f^{\pi}(1,X)-f^{\pi}(0,X))^{2}\right]\right|$
$\leq$ $\sup_{\pi\in\Pi}\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{f}^{\pi}(1,X_{i})-f^{\pi}(1,X_{i})\right)^{2}+\sup_{\pi\in\Pi}\frac{1}{n}\sum_{i=1}^{n}\left(f^{\pi}(0,X_{i})-\widehat{f}^{\pi}(0,X_{i})\right)^{2}$
$+$ $\sup_{\pi\in\Pi}2\left|\frac{1}{n}\sum_{i=1}^{n}\left(f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i})\right)\left(f^{\pi}(0,X_{i})-\widehat{f}^{\pi}(0,X_{i})\right)\right|$
$+$ $\sup_{\pi\in\Pi}2\left|\frac{1}{n}\sum_{i=1}^{n}\left(f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i})\right)\left(f^{\pi}(1,X_{i})-\widehat{f}^{\pi}(1,X_{i})\right)\right|$
$+$ $\sup_{\pi\in\Pi}2\left|\frac{1}{n}\sum_{i=1}^{n}\left(f^{\pi}(0,X_{i})-\widehat{f}^{\pi}(0,X_{i})\right)\left(f^{\pi}(1,X_{i})-\widehat{f}^{\pi}(1,X_{i})\right)\right|$
$+$ $\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum_{i=1}^{n}\left(f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i})\right)^{2}-\mathbb{E}\left[(f^{\pi}(1,X)-f^{\pi}(0,X))^{2}\right]\right|.$
In the above 6 terms, the first five terms are relatively easy to deal with, since their upper bounds can be shown not depend on $\pi$; while the last term
$\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum_{i=1}^{n}\left(f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i})\right)^{2}-\mathbb{E}\left[(f^{\pi}(1,X)-f^{\pi}(0,X))^{2}\right]\right|$
needs to be treated using empirical process theory.
We next bound the first term $\sup_{\pi\in\Pi}\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{f}^{\pi}(1,X_{i})-f^{\pi}(1,X_{i})\right)^{2}$, and the second term can be similarly bounded.
$\sup_{\pi\in\Pi}\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{f}^{\pi}(1,X_{i})-f^{\pi}(1,X_{i})\right)^{2}$
$=$ $\sup_{\pi\in\Pi}\frac{1}{n}\sum_{i=1}^{n}\left(\sum_{a}\pi(a|1,X_{i})\left(\widehat{r}(1,X_{i},a)-r(1,X_{i},a)\right)\right)^{2}$
$\leq$ $\sup_{\pi\in\Pi}\frac{1}{n}\sum_{i=1}^{n}\sum_{a}\pi(a|1,X_{i})\left(\widehat{r}(1,X_{i},a)-r(1,X_{i},a)\right)^{2}\quad\text{(Jensen's Inequality)}$
$\leq$ $\sup_{\pi\in\Pi}\frac{1}{n}\sum_{a}\sum_{i=1}^{n}\pi(a|1,X_{i})\left(\widehat{r}(1,X_{i},a)-r(1,X_{i},a)\right)^{2}$
$\leq$ $\max_{a}\frac{1}{n}\sum_{i=1}^{n}\left[\widehat{r}(1,X_{i},a)-r(1,X_{i},a)\right]^{2}=O_{p}(n^{-\gamma_{2}}),$
where the last equality follows from Assumption 4.4.
We next upper bound the third term. By Cauchy-Schwarz inequality,
$\sup_{\pi\in\Pi}2\left|\frac{1}{n}\sum_{i=1}^{n}\left(f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i})\right)\left(f^{\pi}(0,X_{i})-\widehat{f}^{\pi}(0,X_{i})\right)\right|$
$\lesssim\;$ $\sup_{\pi\in\Pi}\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left[f^{\pi}(0,X_{i})-\widehat{f}^{\pi}(0,X_{i})\right]^{2}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left[f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i})\right]^{2}}$
$\lesssim\;$ $\sqrt{\sup_{\pi\in\Pi}\frac{1}{n}\sum_{i=1}^{n}\left[f^{\pi}(0,X_{i})-\widehat{f}^{\pi}(0,X_{i})\right]^{2}},$
where the last inequality is due to the bounded reward assumption. Note that the above upper bound is simply the square root of the second term, hence is $O_{p}(n^{-\gamma_{2}/2})$.We can bound the 4th and 5th terms similarly. We omit the details to save the space.
It now remains to bound the last term. Define
$\widetilde{\Delta}_{2}(\pi)\equiv\frac{1}{n}\sum_{i=1}^{n}(f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i}))^{2},$
i.e., the oracle estimator of $\Delta_{2}(\pi)$, and let $\Delta_{2i}=(f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i}))^{2}$. Applying the symmetrization technique (Van Der Vaart and Wellner, 1996), we obtain
$\mathbb{E}\sup_{\pi\in\Pi}|\Delta_{2}(\pi)-\widetilde{\Delta}_{2}(\pi)|\leq 2\mathbb{E}\sup_{\pi\in\Pi}\left\lvert\cfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\Delta_{2i}(\pi)\right\rvert$
$=$ $2\mathbb{E}\sup_{\pi\in\Pi}\cfrac{1}{n}\left\lvert\sum_{i=1}^{n}\epsilon_{i}(f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i}))^{2}\right\rvert$
$=$ $2\mathbb{E}\sup_{\pi\in\Pi}\cfrac{1}{n}\left\lvert\sum_{i=1}^{n}\epsilon_{i}\left[\sum_{a}r(1,X_{i},a)\pi(a|1,X_{i})-\sum_{a}r(0,X_{i},a)\pi(a|0,X_{i})\right]^{2}\right\rvert$
$\leq$ $2\underbrace{\mathbb{E}\sup_{\pi\in\Pi}\cfrac{1}{n}\left\lvert\sum_{i=1}^{n}\epsilon_{i}\left[\sum_{a}r(1,X_{i},a)\pi(a|1,X_{i})\right]^{2}\right\rvert}_{E_{1}}+2\underbrace{\mathbb{E}\sup_{\pi\in\Pi}\cfrac{1}{n}\left\lvert\sum_{i=1}^{n}\epsilon_{i}\left[\sum_{a}r(1,X_{i},a)\pi(a|1,X_{i})\right]^{2}\right\rvert}_{E_{2}}$
$+4\underbrace{\mathbb{E}\sup_{\pi\in\Pi}\cfrac{1}{n}\left\lvert\sum_{i=1}^{n}\epsilon_{i}\sum_{a}r(1,X_{i},a)\pi(a|1,X_{i})\sum_{a}r(1,X_{i},a)\pi(a|1,X_{i})\right\rvert}_{E_{3}}.$
It follows that
$E_{1}=\mathbb{E}\sup_{\pi\in\Pi}\cfrac{1}{n}\left\lvert\sum_{i=1}^{n}\epsilon_{i}\left[\sum_{a}r(1,X_{i},a)\pi(a|1,X_{i})\right]^{2}\right\rvert$
$=\;$ $\mathbb{E}\sup_{\pi\in\Pi}\cfrac{1}{n}\left\lvert\sum_{i=1}^{n}\epsilon_{i}\left[r(1,X_{i},1)\pi(1|1,X_{i})\right]^{2}\right\rvert+\mathbb{E}\sup_{\pi\in\Pi}\cfrac{1}{n}\left\lvert\sum_{i=1}^{n}\epsilon_{i}\left[r(1,X_{i},0)\pi(0|1,X_{i})\right]^{2}\right\rvert$
$+\mathbb{E}\sup_{\pi\in\Pi}\cfrac{1}{n}\left\lvert\sum_{i=1}^{n}\epsilon_{i}r(1,X_{i},0)\pi(0|1,X_{i})r(1,X_{i},1)\pi(1|1,X_{i})\right\rvert$.
By the bounded reward and finite VC subgraph dimension assumption (Assumption 4.3) and arguing similarly as in the proof of Lemma 3, the above three terms all are order of $O(\sqrt{\mbox{VC}(\Pi)/n})$. The terms $E_{2}$ and $E_{3}$ can be proved analogously. Applying the bounded difference inequality as in the proof of Lemma 3, we have proved that
$\sup_{\pi\in\Pi}|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)|=O_{p}\left(\sqrt{\mbox{VC}(\Pi)/n}+n^{-\gamma_{2}/2}+\sqrt{\log n/n}\right)$
for the equal opportunity fairness notion. ∎
9.7 Proof of Theorem 3: Counterfactual Fairness
By Theorems 1 and 2, it remains to
  1. 1.
    Establish an upper bound for $\sup_{\pi\in\Pi}\big|\widehat{V}(\pi)-V(\pi)\big|$;
  2. 2.
    Establish an upper bound for
    $\sup_{\pi\in\Pi}\big|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)\big|+\sup_{\pi\in\Pi}\big|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)\big|,$
    under the counterfactual fairness notion.
We present the following lemmas.
Lemma 5 (Counterfactual Action Fairness Evaluation).
For the equal opportunity fairness notion, the following holds:
$\sup_{\pi\in\Pi}|\widehat{\Delta}_{1}(\pi)-\Delta_{1}(\pi)|=O_{p}\left(\sqrt{\mbox{VC}(\Pi)/n}+\sqrt{\log n/n}+\sqrt{d/n}\right).$
Proof.
By the definition, we have
$\sup_{\pi\in\Pi}|\Delta_{1}(\pi)-\widehat{\Delta}_{1}(\pi)|=\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum_{i=1}^{n}\left[\pi(S_{i},X_{i})-\pi(S_{i}^{\prime},\widehat{X}_{i}^{\prime})\right]^{2}-\mathbb{E}\left[\pi(S,X)-\pi(S^{\prime},X^{\prime})\right]^{2}\right|$
$\leq$ $\;\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum_{i=1}^{n}\left[\pi(S_{i},X_{i})-\pi(S_{i}^{\prime},\widehat{X}_{i}^{\prime})\right]^{2}-\mathbb{E}\left[\pi(S,X)-\pi(S^{\prime},\widehat{X}^{\prime})\right]^{2}\right|$
$+\sup_{\pi\in\Pi}\left|\mathbb{E}\left[\pi(S,X)-\pi(S^{\prime},\widehat{X}^{\prime})\right]^{2}-\mathbb{E}\left[\pi(S,X)-\pi(S^{\prime},X^{\prime})\right]^{2}\right|$
$\leq$ $\;\underbrace{\sup_{\begin{subarray}{c}\pi\in\Pi\\ t:\,\lVert t\rVert_{2}\leq\delta\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\left[\pi(S_{i},X_{i})-\pi(S_{i}^{\prime},X_{i}^{\prime}+t)\right]^{2}-\mathbb{E}\left[\pi(S,X)-\pi(S^{\prime},X^{\prime}+t)\right]^{2}\right|}_{I_{1}}$
$+\underbrace{\sup_{\begin{subarray}{c}\pi\in\Pi\\ t:\,\lVert t\rVert_{2}\leq\delta\end{subarray}}\left|\mathbb{E}\left[\pi(S,X)-\pi(S^{\prime},X^{\prime}+t)\right]^{2}-\mathbb{E}\left[\pi(S,X)-\pi(S^{\prime},X^{\prime})\right]^{2}\right|}_{I_{2}},$
with high probability $\delta\leq c\sqrt{d/n}$, since $\lVert\widehat{X}^{\prime}-X^{\prime}\rVert_{2}=O_{p}(\sqrt{d/n})$. This follows directly from Assumption 4.5, allowing $\widehat{\theta}(s)$ for $s=0,1$ to be estimated using the sample average, i.e., $\widehat{\theta}(s)=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}(S_{i}=s)X_{i}$.
The second term $I_{2}=O_{p}(\sqrt{d/n})$ by Assumption 4.6. It remains to deal with the empirical process term $I_{1}.$ Define the function class
$\Pi_{t}=\{\pi(a|s,x+t):\pi\in\Pi,\;t\leq\delta\}.$
Since $\Pi$ is a VC class, and $t$ ranges over a bounded set, the $\varepsilon$-covering number of the function class is upper bounded by
$C\left(\frac{1}{\varepsilon}\right)^{{\mbox{VC}(\Pi)}}\left(\frac{\delta}{\varepsilon}\right)^{d},$
and the corresponding entropy is upper bounded by
$d\log\delta+d\log\left(\frac{1}{\varepsilon}\right)+\mbox{VC}(\Pi)\log\left(\frac{1}{\varepsilon}\right),$
up to a constant. Thus, the transformed class $\{\pi^{2}(a|s,x+t):\pi\in\Pi,\;t\leq\delta\}$ also has bounded entropy (see the argument in the proof of Lemma 3). Therefore, using symmetrization, Dudley’s entropy bound, and bounded difference inequality yields
$I_{1}=O_{p}\left(\sqrt{\mbox{VC}(\Pi)/n}+\sqrt{\log n/n}+\sqrt{d/n}\right).$
This completes the proof. ∎
Lemma 6 (Counterfactual Outcome fairness Evaluation).
For the counterfactual fairness notion, the following holds:
$\sup_{\pi\in\Pi}|\Delta_{2}(\pi)-\widehat{\Delta}_{2}(\pi)|=O_{p}\left(\sqrt{\mbox{VC}(\Pi)/n}+n^{-\gamma_{2}/2}+\sqrt{\log n/n}+\sqrt{d/n}\right).$
Proof.
The proof can be established in a manner similar to the proofs of Lemmas 4 and 5. For brevity, the details are omitted. ∎

### v2

@@
-Double Fairness Policy Learning: Integrating Action Fairness and Outcome Fairness in Decision-making
+
 Abstract
-Fairness is a central pillar of trustworthy machine learning, especially in domains where accuracy- or profit-driven optimization is insufficient.
-While most fairness research focuses on supervised learning, fairness in policy learning remains less explored.
-Because policy learning is interventional, it induces two distinct fairness targets: action fairness (equitable action assignments) and outcome fairness (equitable downstream consequences).
-Crucially, equalizing actions does not generally equalize outcomes when groups face different constraints or respond differently to the same action.
-We propose a novel double fairness learning (DFL) framework that explicitly manages the trade-off among three objectives: action fairness, outcome fairness, and value maximization.
-We integrate fairness directly into a multi-objective optimization problem for policy learning and employ a lexicographic weighted Tchebyshev method that recovers Pareto solutions beyond convex settings, with theoretical guarantees on the regret bounds.
+Fairness has recently received significant attention in machine learning, driven by ethical concerns surrounding algorithmic applications.
+However, fairness in policy learning, where the goal is to make fair decision while simultaneously maximizing expected rewards, has been less explored.
+Most existing works in fairness for policy learning focus primarily on action fairness.
+However, many applications require fairness in both the actions taken and their resulting outcomes.
+This paper tackles the challenge of double fairness in policy learning by considering the trade-off of three objectives simultaneously: action fairness, outcome fairness, and maximizing the value function.
+We propose a novel framework that integrates fairness directly into a multi-objective optimization problem for policy learning.
 Our framework is flexible and accommodates various commonly used fairness notions.
-Extensive simulations demonstrate improved performance relative to competing methods.
-In applications to a motor third-party liability insurance dataset and an entrepreneurship training dataset, DFL substantially improves both action and outcome fairness while incurring only a modest reduction in overall value.
-Keywords: Fairness; Trustworthiness; Policy Learning; Multi-objective Optimization.
+We establish theoretical guarantees for the proposed method and demonstrate its empirical superiority over competing approaches.
+Finally, we apply our method to a motor third-party liability insurance dataset and an entrepreneurship training dataset.
+Our proposed method demonstrates greater fairness on both fairness metrics while having only a modest reduction in overall revenue.
+keywords:
+Fairness; Trustworthiness; Policy Learning; Multi-objective Optimization
 1 Introduction
-As a core aspect of trustworthiness, fairness has become central to machine learning (Dwork et al., 2012; Zemel et al., 2013; Grgic-Hlaca et al., 2016; Hardt et al., 2016; Kusner et al., 2017; Wu et al., 2019; Tan et al., 2022; Chen et al., 2023), motivated by ethical concerns about the societal impacts of algorithmic decision-making.
+As a core aspect of trustworthiness, fairness has become central to machine learning (dwork2012fairness; zemel2013learning; grgic2016case; hardt2016equality; kusner2017counterfactual; wu2019counterfactual; tan2022rise; chen2023learning), motivated by ethical concerns about the societal impacts of algorithmic decision-making.
 In many domains such as healthcare, insurance, lending, and public programs, algorithmic systems can affect large populations at scale.
@@
 Most existing work on algorithmic fairness has focused on supervised learning.
-In contrast, fairness in policy learning, where the goal is to maximize expected reward or social welfare, has received comparatively limited attention (Nabi et al., 2019; Tan et al., 2022; Viviano and Bradic, 2023; Fang et al., 2023; Wang et al., 2025).
+In contrast, fairness in policy learning, where the goal is to maximize expected reward or social welfare, has received comparatively limited attention (nabi2019learning; tan2022rise; viviano2023fair; fang2023fairness; wang2025counterfactually).
 A key distinction is that policy learning is inherently interventional: a policy selects an action $A$, and that action alters subsequent outcomes.
@@
 Importantly, equalizing actions does not, by itself, equalize outcomes when groups face different constraints, environments, or downstream responses to the same action.
-Consistent with this distinction, our analysis of a Belgian motor liability insurance dataset (Dutang and Charpentier, 2019) show that standard policy optimization methods can induce substantial disparities in both action and outcome fairness (Table 1 in Section 6).
+Consistent with this distinction, our analysis of a Belgian motor liability insurance dataset (dutang2019casdatasets) show that standard policy optimization methods can induce substantial disparities in both action and outcome fairness (Table 1 in Section 6).
 This “action–outcome” gap arises naturally in practice.
@@
 Machine learning algorithms are increasingly used to automate risk assessment, pricing, and claims-related decisions, making trustworthy decision rules especially critical.
-For example, UnitedHealth has used the nH Predict platform (Talia, 2024) to assess patient risk and forecast outcomes in ways that inform coverage and pricing.
+For example, UnitedHealth has used the nH Predict platform (talia2024algorithms) to assess patient risk and forecast outcomes in ways that inform coverage and pricing.
 While such tools can improve efficiency, they also raise concerns about algorithmic bias and discrimination against vulnerable populations.
@@
 Hence, a principled framework is needed to characterize and manage these trade-offs, for which we adopt the concept of Pareto optimality.
-Third, computing the relevant Pareto fairness set is nontrivial in practice: standard linear scalarization approaches can miss Pareto-optimal solutions when the policy class is non-convex, as is typical for many policy parameterizations (Schulman et al., 2015).
+Third, computing the relevant Pareto fairness set is nontrivial in practice: standard linear scalarization approaches can miss Pareto-optimal solutions when the policy class is non-convex, as is typical for many policy parameterizations (schulman2015trust).
 We address these challenges with a unified methodology and theory.
 We first provide a systematic analysis of when outcome-fair and doubly fair policies can be achieved, yielding interpretable conditions for feasibility.
-To manage trade-offs, we cast DFL as a multi-objective optimization problem (Miettinen, 1999) and target policies that are Pareto-efficient with respect to the two fairness criteria while achieving high value function.
-To compute Pareto solutions, we adopt a lexicographic weighted Tchebyshev scalarization strategy (Steuer and Choo, 1983), which can recover Pareto solutions without requiring convexity (see Figure 7 in Appendix).
+To manage trade-offs, we cast DFL as a multi-objective optimization problem (miettinen1999nonlinear) and target policies that are Pareto-efficient with respect to the two fairness criteria while achieving high value function.
+To compute Pareto solutions, we adopt a lexicographic weighted Tchebyshev scalarization strategy (steuer1983interactive), which can recover Pareto solutions without requiring convexity (see Figure 7 in Appendix).
 We then develop an estimator that embeds fairness objectives directly into the policy optimization procedure, establish regret bounds that provide rigorous performance guarantees under double fairness constraints.
@@
 Policy learning seeks to maximize expected reward by learning an optimal policy that selects the best action for each situation.
-It has broad applications in precision medicine (Murphy, 2003; Robins, 2004; Qian and Murphy, 2011; Zhao et al., 2012; Wallace and Moodie, 2015; Wang et al., 2018; Shi et al., 2018; Qi et al., 2020; Cui and Tchetgen Tchetgen, 2021; Bian et al., 2023), economics (Manski, 2004; Kitagawa and Tetenov, 2018; Athey and Wager, 2021), and business strategy (Den Boer, 2015; Javanmard and Nazerzadeh, 2019; Bian et al., 2024).
+It has broad applications in precision medicine (Murphy; gest; qian2011performance; zhao2012estimating; wallace2015doubly; wang2018quantile; shi2018high; qi2020multi; cui2021semiparametric; bian2023variable), economics (manski2004statistical; kitagawa2018should; athey2021policy), and business strategy (den2015dynamic; javanmard2019dynamic; bian2024tale).
 This literature largely focuses on value maximization, whereas our work augments policy learning with two complementary fairness targets—action fairness (fairness of decisions) and outcome fairness (fairness of the induced consequences)—and studies how to optimize value under both.
 Work on fairness-aware policy learning is comparatively scarce and typically enforces a single fairness notion.
-Nabi et al.
-(2019) formulate fair policy learning via causal mediation analysis, specifying impermissible causal pathways from the sensitive attribute and optimizing policies under the resulting causal constraints.
-Tan et al.
-(2022), Fang et al.
-(2023), and Cui and Han (2025) develop quantile-optimal decision rules designed to protect disadvantaged groups by controlling tail outcomes.
-Viviano and Bradic (2023) adopt a “first do no harm” principle by characterizing a welfare Pareto frontier and selecting the fairest allocation within that welfare-efficient set, with computation based on mixed-integer optimization.
-Wang et al.
-(2025) study counterfactual fairness in sequential decision making, proposing approaches tailored to reinforcement learning settings.
+nabi2019learning formulate fair policy learning via causal mediation analysis, specifying impermissible causal pathways from the sensitive attribute and optimizing policies under the resulting causal constraints.
+tan2022rise, fang2023fairness, and cui2025policy develop quantile-optimal decision rules designed to protect disadvantaged groups by controlling tail outcomes.
+viviano2023fair adopt a “first do no harm” principle by characterizing a welfare Pareto frontier and selecting the fairest allocation within that welfare-efficient set, with computation based on mixed-integer optimization.
+wang2025counterfactually study counterfactual fairness in sequential decision making, proposing approaches tailored to reinforcement learning settings.
 In contrast to these works, we explicitly distinguish action fairness from outcome fairness and address the double fairness problem with different theoretical tools.
@@
 We also accommodate the degenerate case where $R^{(1)}=R^{(2)}$: for instance, in the government job-training data (Section 6.2), both components correspond to subsequent entrepreneurial activity.
-In dynamic pricing, $R^{(1)}$ typically represents net profit, whereas $R^{(2)}$ might denote demand or consumer surplus—metrics for which fairness has been extensively studied (see, e.g., Cohen et al., 2022; Bian et al., 2026).
+In dynamic pricing, $R^{(1)}$ typically represents net profit, whereas $R^{(2)}$ might denote demand or consumer surplus—metrics for which fairness has been extensively studied (see, e.g., cohen2022price; bian2026beyond).
 Furthermore, in healthcare, $R^{(1)}$ may represent clinical outcomes, while $R^{(2)}$ denotes the allocation of medical resources.
@@
 where $\widehat{r}(s,x,a)$ is the regression function estimator of the reward function $r(s,x,a)$.
-Such a regression approach has been widely used in the policy learning literature; see, e.g., Qian and Murphy (2011); Moodie et al.
-(2012).
+Such a regression approach has been widely used in the policy learning literature; see, e.g., qian2011performance; moodie2012q.
 Then in the value maximization step, the estimated optimal policy is derived by maximizing this estimated value function, within the pre-specified policy class $\Pi$:
@@
 2.2 Fairness Notions
-Recently, many fairness notions have been proposed under supervised learning setting, with particular attention to equal opportunity (Hardt et al., 2016) and counterfactual fairness (Kusner et al., 2017).
+Recently, many fairness notions have been proposed under supervised learning setting, with particular attention to equal opportunity (hardt2016equality) and counterfactual fairness (kusner2017counterfactual).
 Here, we extend these two notions to the policy learning framework and formally define both action fairness and outcome fairness under each concept.
@@
 Equal opportunity requires that individuals with the same observed information $X$ be treated equally across different sensitive groups.
-This criterion is intuitively appealing and has been widely used in practice (Hardt et al., 2016).
-However, because it conditions only on observed covariates, it may fail to account for historical or unobserved sources of bias (Chen et al., 2023).
-To address this, Kusner et al.
-(2017) proposed counterfactual fairness, which adopts a causal perspective by asking whether the decision (or the induced outcome) would have remained unchanged had the individual belonged to a different sensitive group.
+This criterion is intuitively appealing and has been widely used in practice (hardt2016equality).
+However, because it conditions only on observed covariates, it may fail to account for historical or unobserved sources of bias (chen2023learning).
+To address this, kusner2017counterfactual proposed counterfactual fairness, which adopts a causal perspective by asking whether the decision (or the induced outcome) would have remained unchanged had the individual belonged to a different sensitive group.
 Definition 2 (Counterfactual Fairness).
@@
 Therefore, it is crucial to understand how to effectively balance multiple objectives and make optimal trade-offs among them.
-The concept of Pareto set (see, e.g., Miettinen, 1999) offers a valuable framework for this purpose, capturing the idea that improvements in one objective cannot be achieved without compromising at least one other.
+The concept of Pareto set (see, e.g., miettinen1999nonlinear) offers a valuable framework for this purpose, capturing the idea that improvements in one objective cannot be achieved without compromising at least one other.
 By leveraging the idea of Pareto set, we can navigate complex decision-making scenarios and strive for policies that optimize overall benefits across diverse objectives.
@@
 To better characterize the Pareto fairness set, we present the following useful proposition.
-Proposition 3 (Lexicographic Weighted Tchebychef Method (Steuer and Choo, 1983)).
+Proposition 3 (Lexicographic Weighted Tchebychef Method (steuer1983interactive)).
 The Pareto fairness set can be computed via
@@
 Remark 2.
-Viviano and Bradic (2023) adopted the linear scalarization approach (specific formulation discussed in Section 5) to characterize the Pareto set.
+viviano2023fair adopted the linear scalarization approach (specific formulation discussed in Section 5) to characterize the Pareto set.
 However, this method cannot identify Pareto fairness policies in the non-convex regions of the Pareto fairness front (as illustrated in Figure 7 in Section 8.1 of the Supplemental Materials).
@@
 Denote the solution as $\widehat{\pi}_{k}^{*}$.
-7:  Compute $k^{*}=\operatorname*{arg\,max}_{1\leq k\leq K}\widehat{V}_{k}^{*}$ .
-8:  return $\widehat{\pi}=\widehat{\pi}_{k^{*}}^{*}$.
+7:  end for
+8:  Compute $k^{*}=\operatorname*{arg\,max}_{1\leq k\leq K}\widehat{V}_{k}^{*}$ .
+9:  return $\widehat{\pi}=\widehat{\pi}_{k^{*}}^{*}$.
+10:end function
 4 Theoretical Results
@@
 Theorem 1 shows the regret consists of two parts: (i) the uniform policy evaluation error over $\Pi_{p}$ and $\Pi$, and (ii) the estimation error of the Pareto fairness set, measured by the Hausdorff distance $H_{\mathrm{TV}}(\Pi_{p},\widehat{\Pi}_{p})$.
-For (i), if the nuisance functions are estimated with sufficient accuracy and $\Pi$ is a VC subgraph class, then the uniform error admits a suitable bound; see, e.g., Kitagawa and Tetenov (2018); Athey and Wager (2021).
+For (i), if the nuisance functions are estimated with sufficient accuracy and $\Pi$ is a VC subgraph class, then the uniform error admits a suitable bound; see, e.g., kitagawa2018should; athey2021policy.
 We now analyze component (ii), starting with the following assumption.
@@
 Assumption 4.2 imposes a separation condition: policies that are at least $\delta$ away from $\Pi_{p}$ in total variation must incur an objective gap of order $\delta$ under either the criterion $\Delta(\cdot)$ or the weighted Tchebycheff criterion $M_{\alpha}(\cdot)$.
-Such well-separated solution conditions are standard in establishing consistency of $M$- and $Z$-estimators; see, e.g., Van der Vaart (2000).
+Such well-separated solution conditions are standard in establishing consistency of $M$- and $Z$-estimators; see, e.g., van2000asymptotic.
 Theorem 2.
@@
 The policy class $\Pi$ has finite VC-subgraph dimension; that is, $\mbox{VC}(\Pi)$ is bounded.
-Assumption 4.3 is commonly imposed in policy learning literature, see, e.g., Kitagawa and Tetenov (2018); Athey and Wager (2021).
+Assumption 4.3 is commonly imposed in policy learning literature, see, e.g., kitagawa2018should; athey2021policy.
 Common examples of classes with finite VC-subgraph dimension include linear decision rules, decision trees, and sigmoid (logistic) score classes.
@@
 If $r$ and $f$ are estimated using correctly specified parametric models, one typically has $\gamma_{1}=\gamma_{2}=1$.
-Alternatively, flexible nonparametric methods such as sieve estimators (Huang, 1998) can be used.
+Alternatively, flexible nonparametric methods such as sieve estimators (huang1998projection) can be used.
 Under standard smoothness and complexity conditions, sieve estimators can achieve the optimal mean-squared rate $O_{p}(n^{-\frac{2\beta}{2\beta+d}})$, where $\beta$ is the smoothness parameter of the reward function, and $d$ is the dimension of $X$.
-Sieve methods are widely used in the policy evaluation and learning literature; see, e.g., Shi et al.
-(2022); Chen and Qi (2022); Bian et al.
-(2025).
+Sieve methods are widely used in the policy evaluation and learning literature; see, e.g., shi2020statistical; chen2022well; bian2025off.
 For the counterfactual fairness notion, we additionally impose the following Assumption 4.5 to ensure that the counterfactual covariates $X_{s,x}(s^{\prime})$ are identifiable from the observed data.
-Similar assumptions are used in Chen et al.
-(2023) and Wang et al.
-(2025) to study counterfactual fairness.
+Similar assumptions are used in chen2023learning and wang2025counterfactually to study counterfactual fairness.
 Assumption 4.5.
@@
 $(ii)\sup_{\begin{subarray}{c}\pi\in\Pi\\ t:\,\lVert X-t\rVert_{2}\leq\delta\end{subarray}}\mathbb{E}\left[\pi(S,X)-\pi(S,t)\right]^{2}\leq m\delta.$
-Assumption 4.6(i) is standard and has also been imposed by Wang et al.
-(2025) for counterfactual fairness policy learning.
+Assumption 4.6(i) is standard and has also been imposed by wang2025counterfactually for counterfactual fairness policy learning.
 Assumption 4.6(ii) requires only Lipschitz continuity in expected mean square, which is mild enough to accommodate deterministic policies (for which pointwise Lipschitz continuity typically fails).
-A closely related condition is used by Shi et al.
-(2021) when analyzing linear deterministic decision rules; see Assumption A5(ii) therein.
+A closely related condition is used by shi2021concordance when analyzing linear deterministic decision rules; see Assumption A5(ii) therein.
 We now present the main regret bounds for equal opportunity and counterfactual fairness.
@@
 Suppose the assumptions of Theorem 3 hold. In addition, assume Assumptions 4.5 and 4.6. Then, under the counterfactual notion of fairness, we have
-$|\mu(\widehat{\pi})|=O_{p}\big(\sqrt{\mbox{VC}(\Pi)/n}+\sqrt{\log n/n}+\sqrt{d/n}\big)++O_{p}\big(n^{-\frac{\gamma_{1}}{2}}+n^{-\frac{\gamma_{2}}{2}}\big).$
+$|\mu(\widehat{\pi})|=O_{p}\big(\sqrt{\mbox{VC}(\Pi)/n}+\sqrt{\log n/n}+\sqrt{d/n}\big)+O_{p}\big(n^{-\frac{\gamma_{1}}{2}}+n^{-\frac{\gamma_{2}}{2}}\big).$
 Theorems 3 and 4 establish the regret bound of the estimated policy under the notions of equal opportunity fairness and counterfactual fairness, respectively.
@@
 For sieve estimation, the optimal number of basis functions satisfies $L^{\mathrm{opt}}\;=\;n^{\frac{d}{d+2\beta}}.$ Thus, one can first select $L^{\text{opt}}$ (e.g., via cross-validation), and then back out $\beta$ and hence $\gamma_{2}$ accordingly.
-The same logic extends to other nonparametric methods (e.g., kernel or local polynomial estimators), whose optimal bandwidth choices also depend on $\beta$; see, for example, Tsybakov (2009).
+The same logic extends to other nonparametric methods (e.g., kernel or local polynomial estimators), whose optimal bandwidth choices also depend on $\beta$; see, for example, Tsybakov2009.
 5 Numerical Studies
@@
 Data generating procedure.
-We consider the following data-generating process, which is similar to that in Section F.2.1 of Wang et al.
-(2025).
+We consider the following data-generating process, which is similar to that in Section F.2.1 of wang2025counterfactually.
 The sensitive variable is given by $S_{i}\sim\mbox{Bernoulli}(0.35)$, and we consider the case in which the covariates are two-dimensional and are generated by $X_{1i}=0.45S_{i}+\varepsilon_{1i}$, and $X_{2i}=0.85S_{i}+\varepsilon_{2i}$, where $\{\varepsilon_{1i}\}_{i}$ and $\{\varepsilon_{2i}\}_{i}$ are independent error terms drawn from a standard Gaussian distribution.
@@
 Competing methods.
-We compare our proposed method with the standard policy learning approach, and the approach for finding Pareto policies in Viviano and Bradic (2023).
+We compare our proposed method with the standard policy learning approach, and the approach for finding Pareto policies in viviano2023fair.
 Specifically, the standard policy learning approach estimates the optimal policy as
 $\operatorname*{arg\,max}_{\pi\in\Pi}\;\frac{1}{n}\sum_{i=1}^{n}\left[\widehat{r}(S_{i},X_{i},1)-\widehat{r}(S_{i},X_{i},0)\right]\pi(S_{i},X_{i}).$
-As for the approach outlined in Viviano and Bradic (2023), it focuses solely on single fairness policy learning tasks, either outcome fairness or action fairness.
+As for the approach outlined in viviano2023fair, it focuses solely on single fairness policy learning tasks, either outcome fairness or action fairness.
 To build on this, we first consider the original method based on the action fairness metric (VB1) and the outcome fairness metric (VB2).
@@
 6.1 Motor Insurance Liability Claims Data
-In this section, we illustrate our proposed method to a Belgian motor third-party liability dataset from the CASdatasets (Dutang and Charpentier, 2019).
-Similar to Xin and Huang (2024), we use gender as the sensitive variable.
-In the offline dataset, we classify customers as low-risk or high-risk based on the bonus-malus (BM) scale, following common practice in the insurance literature, where drivers are often categorized as good or bad based on their BM score (Frangos and Vrontos, 2001).
+In this section, we illustrate our proposed method to a Belgian motor third-party liability dataset from the CASdatasets (dutang2019casdatasets).
+Similar to xin2024antidiscrimination, we use gender as the sensitive variable.
+In the offline dataset, we classify customers as low-risk or high-risk based on the bonus-malus (BM) scale, following common practice in the insurance literature, where drivers are often categorized as good or bad based on their BM score (frangos2001design).
 The BM scale ranges from 0 to 22, with higher values indicating greater risk.
@@
 The non-sensitive covariates include the number of claims, policyholder age, coverage plan, if the vehicle is part of a fleet, and the coverage duration.
-Since the dataset does not contain premium information, we generate the pure premium amount using the frequency-severity approach, as outlined by Baumann and Loi (2023).
+Since the dataset does not contain premium information, we generate the pure premium amount using the frequency-severity approach, as outlined by baumann2023fairness.
 This approach sets pure premiums based on the predicted risk.
@@
 6.2 Entrepreneurship Training Data
-In this subsection, we illustrate our proposed method to the entrepreneurship training data for North America undergraduate students from Lyons and Zhang (2017).
+In this subsection, we illustrate our proposed method to the entrepreneurship training data for North America undergraduate students from lyons2017impact.
 This study examined whether entrepreneurship programs have varying impacts on certain minority groups, such as females or non-Caucasians.
@@
 In this subsection, we show that the limitation of the linear scalarization approach via the following simple example.
-(a) Linear scalarization.
-(b) Weighted Tchebycheff scalarization.
-Figure 7: Linear scalarization approach cannot identify Pareto polices in the non-convex regions of the Pareto front, as these solutions lack supporting hyperplanes (Miettinen, 1999).
+((a)) Linear scalarization.
+((b)) Weighted Tchebycheff scalarization.
+Figure 7: Linear scalarization approach cannot identify Pareto polices in the non-convex regions of the Pareto front, as these solutions lack supporting hyperplanes (miettinen1999nonlinear).
 In contrast, weighted Tchebycheff scalarization can find all Pareto policies, regardless of the shape of the objective space.
-Figure adapted from Lin et al.
-(2024).
+Figure adapted from lin2024smooth.
 Consider the following data generation procedure: $S\sim\mbox{Ber}(0.5)$, $X\sim\mbox{Ber}(0.6)$, R=S+X+0.4A0.5ASR=S+X+0.4A-0.5AS.
@@
 8.4 Further Details on Premium in the Insurance Data
-In the offline dataset, we classify customers as low-risk or high-risk based on the BM scale, following common practice in the insurance literature, where drivers are often categorized as good or bad based on their BM score (Frangos and Vrontos, 2001).
+In the offline dataset, we classify customers as low-risk or high-risk based on the BM scale, following common practice in the insurance literature, where drivers are often categorized as good or bad based on their BM score (frangos2001design).
 The BM scale ranges from 0 to 22, with higher values indicating greater risk.
 We define customers with BM greater than five as high-risk and the rest as low-risk.
-Since the dataset does not contain premium information, we generate the pure premium amount using the frequency-severity approach, as outlined by Baumann and Loi (2023).
+Since the dataset does not contain premium information, we generate the pure premium amount using the frequency-severity approach, as outlined by baumann2023fairness.
 This approach sets pure premiums based on the predicted risk.
@@
 Proof.
-To show all of the elements in $\Pi_{p}$ is Pareto optimal solution, see Theorem 5.4.1 in Miettinen (1999); while the other direction can be proved using Theorem 5.4.2 in Miettinen (1999). ∎
+To show all of the elements in $\Pi_{p}$ is Pareto optimal solution, see Theorem 5.4.1 in miettinen1999nonlinear; while the other direction can be proved using Theorem 5.4.2 in miettinen1999nonlinear.
+∎
 9.4 Proof of Theorem 1
@@
 $\sqrt{\max_{a}\frac{1}{n}\sum_{i=1}^{n}(\widehat{r}(S_{i},X_{i},a)-r(S_{i},X_{i},a))^{2}}=O_{p}\left(n^{-\frac{\gamma_{1}}{2}}\right),$
-by Assumption 4.4. Using standard empirical process techniques (see, e.g., Van Der Vaart and Wellner (1996)), the second term is of order
+by Assumption 4.4. Using standard empirical process techniques (see, e.g., van1996weak), the second term is of order
 $O_{p}\left(\sqrt{\frac{\mathrm{VC}(\Pi)}{n}}+\sqrt{\frac{\log n}{n}}\right).$
@@
 $\{\pi^{2}(1,\cdot):\pi\in\Pi\},\qquad\{\pi^{2}(0,\cdot):\pi\in\Pi\},\qquad\{\pi(1,\cdot)\pi(0,\cdot):\pi\in\Pi\}$
-are still VC-type, because squaring is a monotone Lipschitz map on $[0,1]$ and products of bounded VC-type classes remain VC-type (see Lemma 2.6.20 in Van Der Vaart and Wellner (1996)).
+are still VC-type, because squaring is a monotone Lipschitz map on $[0,1]$ and products of bounded VC-type classes remain VC-type (see Lemma 2.6.20 in van1996weak).
 Thus, a direct application of Dudley’s entropy integral bound yields that
@@
 $\widetilde{\Delta}_{2}(\pi)\equiv\frac{1}{n}\sum_{i=1}^{n}(f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i}))^{2},$
-i.e., the oracle estimator of $\Delta_{2}(\pi)$, and let $\Delta_{2i}=(f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i}))^{2}$. Applying the symmetrization technique (Van Der Vaart and Wellner, 1996), we obtain
+i.e., the oracle estimator of $\Delta_{2}(\pi)$, and let $\Delta_{2i}=(f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i}))^{2}$. Applying the symmetrization technique (van1996weak), we obtain
 $\mathbb{E}\sup_{\pi\in\Pi}|\Delta_{2}(\pi)-\widetilde{\Delta}_{2}(\pi)|\leq 2\mathbb{E}\sup_{\pi\in\Pi}\left\lvert\cfrac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\Delta_{2i}(\pi)\right\rvert$